### v4

# JaxMARL: Multi-Agent RL Environments and Algorithms in JAX

Abstract.
Benchmarks play an important role in the development of machine learning algorithms. For example, research in reinforcement learning (RL) has been heavily influenced by available environments and benchmarks. However, RL environments are traditionally run on the CPU, limiting their scalability with typical academic compute. Recent advancements in JAX have enabled the wider use of hardware acceleration to overcome these computational hurdles, enabling massively parallel RL training pipelines and environments. This is particularly useful for multi-agent reinforcement learning (MARL) research. First of all, multiple agents must be considered at each environment step, adding computational burden, and secondly, the sample complexity is increased due to non-stationarity, decentralised partial observability, or other MARL challenges. In this paper, we present JaxMARL, the first open-source code base that combines ease-of-use with GPU enabled efficiency, and supports a large number of commonly used MARL environments as well as popular baseline algorithms. When considering wall clock time, our experiments show that per-run our JAX-based training pipeline is up to 12500x faster than existing approaches. This enables efficient and thorough evaluations, with the potential to alleviate the evaluation crisis of the field. We also introduce and benchmark SMAX, a vectorised, simplified version of the popular StarCraft Multi-Agent Challenge, which removes the need to run the StarCraft II game engine. This not only enables GPU acceleration, but also provides a more flexible MARL environment, unlocking the potential for self-play, meta-learning, and other future applications in MARL. We provide code at https://github.com/flairox/jaxmarl.
Key words and phrases:
Multi-Agent Reinforcement Learning, JAX, Benchmarks1. Introduction
Benchmarks play a pivotal role in the development of new single and multi-agent reinforcement learning (MARL) algorithms by defining problems, enabling comparisons, and focusing efforts. For example, in recent years, Go and Chess drove the development of MuZero Schrittwieser et al. (2020) while decentralised StarCraft Micromanagement Foerster et al. (2017) and later the StarCraft Multi-Agent Challenge (SMAC, Samvelyan et al., 2019) resulted in the development of algorithms such as QMIX Rashid et al. (2020), a popular MARL technique.
Data transfer between the CPU (where the environment is simulated) and the GPU (where the agents are evaluated) is a crucial bottleneck for simulation speed. Simulation speed in turn is vital for progress in reinforcement learning (RL) because RL algorithms often require a large number of environment interactions. This problem is even worse in MARL, where non-stationarity and decentralised partial observability greatly worsen the sample complexity Bernstein et al. (2002). Hardware acceleration and parallelisation are crucial to alleviating this, but current acceleration and parallelisation methods are typically not implemented in Python, reducing their accessibility for most machine learning researchers Shacklett et al. (2023); Weng et al. (2022). For example, the extremely efficient Hanabi library Hu and Foerster (2020) from Meta-AI research is implemented in C++ and has seen relatively little adoption by the community. However, recent advances in JAX Bradbury et al. (2018) have opened up new possibilities for using Python code directly with hardware accelerators, enabling the wider use of massively parallel RL training pipelines and environments.
The JAX Bradbury et al. (2018) library provides composable function transformations, allowing for automatic vectorisation, device parallelisation, automatic differentiation and just-in-time (JIT) compilation with XLA Sabne (2020), for device-agnostic optimisation. Using JAX, both the environment rollouts and model training can happen on a hardware accelerator (such as a GPU or TPU), removing the cost of data transfer between devices and allowing for significant parallelisation. Recently, PureJaxRL Lu et al. (2022a, 2023b) has demonstrated the power of this end-to-end JAX-based approach; running both the environment and the model training on a GPU yields a 4000x speedup over a “traditional” pipeline with a GPU-trained policy but a CPU-based environment.
These accelerations could substantially advance RL and MARL research by quickening the testing and iteration of ideas. Furthermore, they lower computational hurdles for in-depth MARL research, enabling researchers to utilise billions of frames and extract more performance from single GPUs.
Alongside the current computational issues faced by MARL researchers, recent work also highlights issues with the evaluation standards and use of benchmarks in the MARL community. In particular, MARL papers typically only test on a few domains. Of the 75 recent MARL papers analysed by Gorsane et al. (2022), 50% used only one evaluation environment and a further 30% used only two. While SMAC and MPE Lowe et al. (2017), the two most used environments, have various tasks or maps, the lack of a standard set raises the risk of biased comparisons and incorrect conclusions. This leads to environment overfitting and unclear progress markers.
Instead, novel MARL methods should be tested on a wide range of domains to accurately evaluate their limits and enable better comparisons. The likely issue preventing this is the lack of a unified codebase and the computational burden of further evaluation.
This paper presents JaxMARL, a Python library that for the first time brings together JAX implementations of eight common MARL environments under one API. We additionally provide JAX implementations for four state-of-the-art algorithms, allowing for end-to-end JAX-based training pipelines in a similar fashion to PureJaxRL. As outlined in Figure 1, we present a library with end-to-end hardware-accelerated training, simple Python implementations, and a broad range of MARL environments. By alleviating computational constraints, JaxMARL allows rapid evaluation of novel methods across a broad set of domains, and hence has the potential to be a powerful tool to address MARL’s evaluation crisis. Specifically, we find that JaxMARL achieves over 12500x speedup compared to “conventional” aproaches.
We also create SMAX, a JAX-based simplification of the centralised training with decentralised execution (CTDE) benchmarks SMAC Samvelyan et al. (2019) and SMACv2 Ellis et al. (2022). SMAX features simplified dynamics, greater flexibility and a more sophisticated but fully-decentralised heuristic AI, while retaining the high-dimensional observation space, complex unit type interactions and procedural scenario generation that lend SMAC and SMACv2 much of their difficulty.
As shown in Figure 2, in addition to SMAX, our library includes the most popular environments from several MARL settings. For centralised training with decentralised execution (CTDE), we include the Multi-Agent Particle Environments (MPE) Lowe et al. (2017), and Multi-Agent Brax (MABrax). Meanwhile, for zero-shot coordination (ZSC) and ad-hoc teamplay, we include Hanabi and Overcooked. Lastly, from the general-sum literature, we include the CoinGame and Spatial-Temporal Representations of Matrix Games (STORM), a representation of matrix games as grid-world scenarios with temporally extended actions. JaxMARL provides the first JAX implementation of these environments and unifies them in a single codebase.
We additionally provide JAX implementations of Independent PPO (IPPO) Schulman et al. (2017); de Witt et al. (2020), QMIX, VDN Sunehag et al. (2017) and Independent -Learning (IQL) Mnih et al. (2015), four of the most common MARL algorithms, allowing new techniques to be easily benchmarked against existing practices. We will extend this list before the camera-ready copy, e.g. with the popular MAPPO Yu et al. (2022) algorithm.
2. Background
2.1. Hardware Accelerated Environments
JAX enables the use of Python code with any hardware accelerator, allowing researchers to write hardware-accelerated code easily. Within the RL community, writing environment code in JAX has gained recent popularity. This brings two chief advantages: firstly, environments written in JAX can be very easily parallelised by using JAX’s vmap operation, which vectorises a function across an input dimension, and secondly writing the environment in JAX allows the agent and environment to be co-located on the GPU, which eliminates the time taken to copy between CPU and GPU memory. Combined, these two factors bring significant increases in training speed, with PureJaxRL Lu et al. (2022a) achieving a x speedup over traditional training in single-agent settings.
2.2. SMAC
StarCraft is a popular environment for testing RL algorithms. It typically features features a centralised controller issuing commands to balance micromanagement, the low-level control of individual units, and macromanagement, the high level plans for economy and resource management.
SMAC Samvelyan et al. (2019), instead, focuses on decentralised unit micromanagement across a range of scenarios divided into three broad categories: symmetric, where each side has the same units, asymmetric, where the enemy team has more units, and micro-trick, which are scenarios designed specifically to feature a particular StarCraft micromanagement strategy. SMACv2 Ellis et al. (2022) demonstrates that open-loop policies can be effective on SMAC and adds additional randomly generated scenarios to rectify SMAC’s lack of stochasticity. However, both of these environments rely on running the full game of StarCraft II, which severely increases their CPU and memory requirements. SMAClite Michalski et al. (2023) attempts to alleviate this computational burden by recreating the SMAC environment primarily in NumPy, with some core components written in C++. While this is much more lightweight than SMAC, it cannot be run on a GPU and therefore cannot be parallelised effectively with typical academic hardware, which commonly has very few CPU cores compared to industry clusters.
3. JaxMARL
We present JaxMARL, a library containing simple and accessible JAX implementations of popular MARL environments and algorithms. JAX enables significant acceleration and parallelisation over existing implementations. To the best of our knowledge, JaxMARLis the first open source library that provides JAX-based implementations of a wide range of MARL environments and baselines.
3.1. API
The interface of JaxMARL is inspired by PettingZoo Terry et al. (2021) and Gymnax. We designed it to be a simple and easy-to-use interface for a wide-range of MARL problems. An example of instantiating an environment from JaxMARL’s registry and executing one transition is presented in Figure 3.
As JAX’s JIT compilation requires pure functions, our step method has two additional inputs compared to PettingZoo’s. The state object stores the environment’s internal state and is updated with each call to step, before being passed to subsequent calls. Meanwhile, key_step is a pseudo-random key, consumed by JAX functions that require stochasticity.
This key is separated from the internal state for clarity.
Similar to PettingZoo, the remaining inputs and outputs are dictionaries keyed by agent names, allowing for differing action and observation spaces.
However, as JAX’s JIT compilation requires arrays to have static shapes, the total number of agents in an environment cannot vary during an episode.
Thus, we do not use PettingZoo’s agent iterator.
Instead, the maximum number of agents is set upon environment instantiation and any agents that terminate before the end of an episode pass dummy actions thereafter.
As asynchronous termination is possible, we signal the end of an episode using a special "__all__" key within done.
The same dummy action approach is taken for environments where agents act asynchronously (e.g. turn-based games).
To ensure clarity and reproducibility, we keep strict registration of environments with suffixed version numbers, for example “MPE Simple Spread V3”. Whenever JaxMARL environments correspond to existing CPU-based implementations, the version numbers match.
3.2. Environments
JaxMARL contains a diverse range of environments, all implemented in JAX. We also introduce SMAX, a SMAC-like environment implemented entirely in JAX. In this section we introduce these environments and provide details on their implementations.
SMAX
The StarCraft Multi-Agent Challenge (SMAC) is a popular benchmark but has a number of shortcomings. First, as noted and addressed in prior work Ellis et al. (2022), it is not sufficiently stochastic to require complex closed-loop policies. Additionally, SMAC relies on StarCraft II as a simulator. While this allows SMAC to use the wide range of units, objects and terrain available in StarCraft II, running an entire instance of StarCraft II is slow Michalski et al. (2023) and memory intensive. StarCraft II runs on the CPU and therefore SMAC’s parallelisation is severely limited with typical academic compute.
| Scenario | Ally Units | Enemy Units | Start Positions |
| 2s3z | 2 stalkers and 3 zealots | 2 stalkers and 3 zealots | Fixed |
| 3s5z | 3 stalkers and 5 zealots | 3 stalkers and 5 zealots | Fixed |
| 5m_vs_6m | 5 marines | 6 marines | Fixed |
| 10m_vs_11m | 10 marines | 11 marines | Fixed |
| 27m_vs_30m | 27 marines | 30 marines | Fixed |
| 3s5z_vs_3s6z | 3 stalkers and 5 zealots | 3 stalkers and 6 zealots | Fixed |
| 3s_vs_5z | 3 stalkers | 5 zealots | Fixed |
| 6h_vs_8z | 6 hydralisks | 8 zealots | Fixed |
| smacv2_5_units | 5 uniformly randomly chosen | 5 uniformly randomly chosen | SMACv2-style |
| smacv2_10_units | 10 uniformly randomly chosen | 10 uniformly randomly chosen | SMACv2-style |
| smacv2_20_units | 20 uniformly randomly chosen | 20 uniformly randomly chosen | SMACv2-style |
Using the StarCraft II game engine constrains environment design. For example, StarCraft II groups units into three races and does not allow units of different races on the same team, limiting the variety of scenarios that can be generated. Secondly, SMAC does not support a competitive self-play setting without significant engineering work. The purpose of SMAX is to address these limitations. It provides access to a SMAC-like, hardware-accelerated, customisable environment that supports self-play and custom unit types.
Units in SMAX are modelled as circles in a two-dimensional continuous space. SMAX makes a number of additional simplifications to the dynamics of StarCraft II, details of which are given in Appendix A.1.
SMAX also features a different, and more sophisticated, heuristic AI. The heuristic in SMAC simply moves to a fixed location Michalski et al. (2023), attacking any enemies it encounters along the way, and the heuristic in SMACv2 globally pursues the nearest agent. Thus the SMAC AI often does not aggressively pursue enemies that run away, and cannot generalise to the SMACv2 start positions, whereas the SMACv2 heuristic AI conditions on global information and is exploitable because of its tendency to flip-flop between two similarly close enemies. SMAC’s heuristic AI must be coded in the map editor, which does not provide a simple coding interface.
In contrast, SMAX features a decentralised heuristic AI that can effectively find enemies without requiring the global information of the SMACv2 heuristic. This guarantees that in principle a 50% win rate is always achievable by copying the decentralised heuristic policy exactly. This means any win-rate below 50% represents a concrete failure to learn.
SMAX scenarios incorporate both a number of the original scenarios from SMAC and scenarios similar to those found in SMACv2. The latter sample units uniformly across all SMAX unit types (stalker, zealot, hydralisk, zergling, marine, marauder) and ensure fairness by having identical team composition for the enemy and ally teams. We provide more details on SMAX in Appendix A.1.
Overcooked
Inspired by the popular videogame of the same name, Overcooked is commonly used for assessing fully cooperative and fully observable Human-AI task performance. The aim is to quickly prepare and deliver soup, which involves putting three onions in a pot, cooking the soup, and serving it into bowls. Two agents, or cooks, must coordinate to effectively divide the tasks to maximise their common reward signal. Our implementation mimics the original from Overcooked-AI (Carroll et al., 2019), including all five original layouts and a simple method for creating additional ones. For a discussion on the limitations of the Overcooked-AI environment, see Lauffer et al. (2023).
Hanabi
Hanabi is a fully cooperative partially observable multiplayer card game, where players can observe other players’ cards but not their own. To win, the team must play a series of cards in a specific order while sharing only a limited amount of information between players. As reasoning about the beliefs and intentions of other agents is central to performance, it is a common benchmark for ZSC and ad-hoc teamplay research. Our implementation is inspired by the Hanabi Learning Environment Bard et al. (2020) and includes custom configurations for varying game settings, such as the number of colours/ranks, number of players, and number of hint tokens. Compared to the Hanabi Learning Environment, which is written in C++ and split over dozens of files, our implementation is a single easy-to-read Python file, which simplifies interfacing with the library and running experiments.
Multi-Agent Particle Environments (MPE)
The multi-agent particle environments feature a 2D world with simple physics where particle agents can move, communicate, and interact with fixed landmarks. Each specific environment varies the format of the world and the agents’ abilities, creating a diverse set of tasks that include both competitive and cooperative settings. We implement all the MPE scenarios featured in the PettingZoo library and the transitions of our implementation map exactly to theirs. We additionally include a fully cooperative predator-prey variant of simple tag, presented in Peng et al. (2021). The code is structured to allow for straightforward extensions, enabling further tasks to be added.
Multi-Agent Brax (MABrax)
MABrax is a derivative of Multi-Agent MuJoCo Peng et al. (2021), an extension of the MuJoCo Gym environment Todorov et al. (2012) that is commonly used for benchmarking continuous multi-agent robotic control.
Our implementation utilises BraxFreeman et al. (2021) as the underlying physics engine and includes five of Multi-Agent MuJoCo’s multi-agent factorisation tasks, where each agent controls a subset of the joints and only observes the local state.
The included tasks, illustrated in Figure 2, are: ant_4x2, halfcheetah_6x1, hopper_3x1, humanoid_9|8, and walker2d_2x3.
The task descriptions mirror those from Gymnasium-Robotics de Lazcano et al. (2023).
| Environment | Original, 1 Env | Jax, 1 Env | Jax, 100 Envs | Jax, 10k Envs | Maximum Speedup |
| MPE Simple Spread | |||||
| MPE Simple Reference | |||||
| Switch Riddle | |||||
| Hanabi | |||||
| Overcooked | |||||
| MABrax Ant 4x2 | |||||
| Starcraft 2s3z | |||||
| Starcraft 27m vs 30m | |||||
| STORM | – | – | |||
| Coin Game |
Coin Game
Coin Game is a two-player grid-world environment which emulates social dilemmas such as the iterated prisoner’s dilemma Snyder (1971). Used as a benchmark for the general-sum setting, it expands on simpler social dilemmas by adding a high-dimensional state. Two players, ‘red’ and ‘blue’ move in a grid world and are each awarded 1 point for collecting any coin. However, ‘red’ loses 2 points if ‘blue’ collects a red coin and vice versa. Thus, if both agents ignore colour when collecting coins their expected reward is 0. Further details are provided in Appendix A.2.
Spatial-Temporal Representations of Matrix Games (STORM)
Inspired by the “in the Matrix” games in Melting Pot 2.0 Agapiou et al. (2022), the STORM (Khan et al., 2022) environment expands on matrix games by representing them as grid-world scenarios. Agents collect resources which define their strategy during interactions and are rewarded based on a pre-specified payoff matrix. This allows for the embedding of fully cooperative, competitive or general-sum games, such as the prisoner’s dilemma Snyder (1971). Thus, STORM can be used for studying paradigms such as opponent shaping, where agents act with the intent to change other agents’ learning dynamics, which has been empirically shown to lead to more prosocial outcomes (Foerster et al., 2018; Khan et al., 2022; Lu et al., 2022b; Zhao et al., 2022). Compared to the Coin Game or matrix games, the grid-world setting presents a variety of new challenges such as partial observability, multi-step agent interactions, temporally-extended actions, and longer time horizons. Unlike the “in the Matrix” games from Melting Pot, STORM features stochasticity, increasing the difficulty Ellis et al. (2022). A further environment specification is provided in Appendix A.3.
Switch Riddle
Originally used to illustrate the Differentiable Inter-Agent Learning algorithm Foerster et al. (2016), Switch Riddle is a simple cooperative communication environment that we include as a debugging tool. prisoners held by a warden can secure their release by collectively ensuring that each has passed through a room with a light bulb and a switch. Each day, a prisoner is chosen at random to enter this room. They have three choices: do nothing, signal to the next prisoner by toggling the light, or inform the warden they think all prisoners have been in the room. The game ends when a prisoner informs the warden or the maximum time steps are reached. The rewards are +1 if the prisoner informs the warden, and all prisoners have been in the room, -1 if the prisoner informs the warden before all prisoners have taken their turn, and 0 otherwise, including when the maximum time steps are reached. We benchmark using the implementation from Zhang et al. (2022).
3.3. Algorithms
In this section, we present our re-implementation of four well known MARL baseline algorithms using JAX. The primary objective of these baselines is to provide a structured framework for developing MARL algorithms leveraging the advantages of the JaxMARL environments. All of the training pipelines are fully compatible with JAX’s JIT and VMAP functions, resulting in a significant acceleration of both the training and metric evaluation processes. This enables parallelisation of training across various seeds and hyperparameters on a single machine in parallel. We follow the CleanRL philosophy of providing clear, single-file implementations Huang et al. (2022).
IPPO
-learning Methods
Our -Learning baselines, including Independent -Learning (IQL) Tampuu et al. (2017), Value Decomposition Networks (VDN) Sunehag et al. (2018), and QMIX Rashid et al. (2018), have been implemented in accordance with the PyMARL codebase Rashid et al. (2018) to ensure consistency with published results and enable direct comparisons with PyTorch. Our baselines natively support aggregating trajectories from batched environments, simplifying parallelisation. This approach is more convenient than managing environments on distinct threads and subsequently aggregating results, as done in PyMARL. We provide a brief overview of the implemented baselines in the Appendix.
4. Results
In our results, we aim to demonstrate the speed and correctness of our environments and algorithms.In several cases, minor changes to the environments mean that our environments do not exactly match the originals on a step-by-step level. We therefore demonstrate the correctness in different ways for each environment and discuss each separately. By combining this evidence, we demonstrate that our library provides overall correct and far quicker baselines on a wide range of sufficiently correct and easily-modifiable environments.
4.1. Environment Speed
We measure the performance of our environments in steps per second when using random actions and compare to the original environments in Table 2 and Figure 4. All results were collected on a single NVIDIA A100 GPU and AMD EPYC 7763 64-core processor. Environments were rolled out for 1000 sequential steps. Many environments have comparable performance to JaxMARL when comparing single environments, but the ease of parallelisation with Jax allows for more efficient scaling compared to CPU-based environments. For example, MPE Simple Spread’s JAX implementation is 2̃0x slower than the original when comparing a single environment, but even when only running environments in parallel, the JAX environment is already over x faster. When considering environments, the JAX versions are much faster, achieving speedups of up to x over the single-threaded environment (in the case of Overcooked). Running this many environments in parallel using CPU environments would require a large CPU cluster and sophisticated communication mechanisms. This engineering is typically beyond the resources of academic labs, and therefore JaxMARL can unlock new research directions for such institutions.
4.2. Algorithm Speed
We investigate the speed of our IPPO implementation in Figure 5. By vectorising over agents, it is possible to train a vast number of agents in a fraction of the time it takes to train a single agent without hardware acceleration. For MPE, it is possible to train 1024 teams in seconds, which is less than seconds per teams of agents. A single run of MARLLIB’s IPPO implementation on the same hardware takes around seconds on average. This represents a speedup of over x.
Our JAX-based -learning algorithms also offer significant speed advantages. In Figure 5(a), training a single IQL, VDN, or QMIX policy in MPE takes seconds while using PyMarl takes over an hour. Training 1024 QMIX learners in a batch requires seconds, which translates to seconds per learner, indicating a x speedup. This speedup is not as large as for IPPO because -learning baselines are typically trained with fewer parallel environments. In our experiments, we used 8 parallel environments for -learning compared to the 25 or 64 used for PPO. This difference is due to -learners benefiting more from a buffer with trajectories collected by different policies, resulting in a more frequent policy update, rather than collecting many trajectories with the same policy in parallel.
For SMAX, we compare our vectorised IPPO baseline to the MAPPO implementation provided in Sun et al. (2023). MAPPO utilises an RNN and IPPO uses a feed forward network. This was run on a machine with a 64-core CPU and NVIDIA 2080Ti GPU. Additionally, as discussed in Section 3.2, SMAC and SMAX are different environments. These caveats aside, the differences in performance are so striking that we believe this clearly demonstrates the advantages of our approach. We trained 512 SMAX teams on 2s3z in under 33 minutes, whereas a single training run of PyTorch IPPO implementation takes 44 hours on average. This is roughly a x speedup.
4.3. Algorithm Correctness
We verify the correctness of our algorithm implementations by comparing to baselines from other libraries on the MPE Simple Spread and Simple Speaker Listener environments. For IPPO we report the mean return across seeds in Figure 4(b). Results were collected on the same hardware as listed in Section 4.1. Our IPPO implementation obtains the same performance as MARLLIB and runs x quicker, taking only ten seconds to train.
For the -learning algorithms, we verify the correctness by comparing with PyMARL implementations of the same algorithms on the MPE Simple Spread and Simple Speaker Listener environments. IQL, VDN and QMIX all obtain the same or better results than their PyMARL counterparts. The returns are from greedy policies and averaged across 8 runs. The hyperparameters used are from the PyMARL library.
4.4. Environment Correctness
MPE
Our MPE environment corresponds exactly to the PettingZoo implementation. We validate this for each environment using a uniform-random policy on rollouts, ensuring all observations and rewards are within a tolerance of at each transition. This tolerance accounts for non-determinism due to running floating point computation on the GPU. The correspondence is also shown through the performance of IPPO in Figure 4(b) and the -learning algorithms in Figures 5(b) and 5(c) respectively, as the performance of these algorithms is inline with existing baselines Yu et al. (2022). We additionally report training performance for IQL on the remaining MPE environments in Appendix C.2.
Overcooked
The transition dynamics of our Overcooked implementation match those of the Overcooked-AI implementation. We demonstrate this by training an IPPO policy on our implementation and evaluating the policy on both our Overcooked implementation and the original at regular intervals. Results are illustrated in Figure 6(a) and performance is similar, demonstrating their equivalence.
SMAX
SMAX and SMAC are different environments. However, we demonstrate some similarity between them by comparing our IPPO and MAPPO implementations against MAPPO results on SMAC, using the implementation from Sun et al. (2023). We show this in Figure 8. SMAX and SMAC have different opponent policies and dynamics, which makes this comparison more qualitative than precise. We describe the differences between the two in more depth in in the supplementary material. However, despite these differences, the environments seem similarly difficult, with some environments being more difficult in SMAC, and some more difficult in SMAX. This is shown in Figure 8 and in the supplementary material.
MABrax
As Brax differs subtly from MuJoCo, MABrax does not correspond to MAMuJoCo but the learning dynamics are qualitatively similar.
To demonstrate this, we report mean training return across 10 seeds for IPPO on ant_4x2 in Figure 6(b), and our results are in line with the performance of TRPO reported in Kuba et al. (2021).
We report the performance of IPPO on HalfCheetah and Walker in Appendix C.1, the results are also in line with TRPO.
| Setting | Recommended Environments |
| CTDE | SMAX (all scenarios), Hanabi (2-5 players), Overcooked |
| Zero-shot Coordination | Hanabi (2 players), Overcooked (5 basic scenarios) |
| General-Sum | STORM (iterated prisoner’s dilemma), STORM (matching pennies) |
| Cooperative Continuous Actions | MABrax |
Hanabi
Our implementation does not correspond exactly to the Hanabi Learning Environment as we use a subtly different observation space, with the reasoning given in Appendix A.4. To demonstrate qualitative similarity, we train IPPO on Hanabi in self-play with 2 players, with the mean test return across 3 seeds reported in Figure 6(c).
STORM, Coin Game & Switch Riddle
STORM differs from Melting Pot 2.0 significantly, making direct comparisons challenging, with differences discussed in Appendix A.3. Furthermore, STORM and Coin Game are general-sum games, so the environment returns of IPPO in self-play would not be a good indicator of performance. Switch Riddle is a simple diagnostic environment – we do not use it for thorough evaluations.
5. Evaluation Recommendations
Previous work Gorsane et al. (2022) has found significant differences in the evaluation protocols between MARL research works. We identify four main research areas that would benefit from our library: cooperative centralised training with decentralised execution (CTDE) Foerster et al. (2016), zero-shot coordination Hu et al. (2020), general-sum games, and cooperative continuous action methods.
To aid comparisons between methods, we recommend standard minimal sets of evaluation environments for each of these settings in Table 3. It’s important to note that these are minimal and we encourage as broad an evaluation as possible. For example, in the zero-shot coordination setting, all methods should be able to evaluate on Hanabi and Overcooked. However, it may also be possible to evaluate such methods on the SMACv2 settings of SMAX. Similarly, SMAX could be used to evaluate two-player zero-sum methods by training in self-play. For some settings, such as continuous action environments and general-sum games, there is only one difficult environment. We encourage further development of JAX-based environments in these settings to improve the quality of evaluation.
6. Related Work
Several open-source libraries exist for both MARL algorithms and environments. The popular library PyMARL Samvelyan et al. (2019) provides PyTorch implementations of QMIX, VDN and IQL and integrates easily with SMAC. E-PyMARL Papoudakis et al. (2021) extends this by adding the actor-critic algorithms MADDPG Lowe et al. (2017), MAA2C Mnih et al. (2016), IA2C Mnih et al. (2016), and MAPPO, and supports the SMAC, Gym Brockman et al. (2016), Robot Warehouse Christianos et al. (2020), Level-Based Foraging Christianos et al. (2020), and MPE environments. Recently released MARLLib Hu et al. (2022) is instead based on the open-source RL library RLLib Liang et al. (2018) and combines a wide range of competitive, cooperative and mixed environments with a broad set of baseline algorithms. Meanwhile, MALib Zhou et al. (2023) focuses on population-based MARL across a wide range of environments. However, none of these frameworks feature hardware-accelerated environments and thus lack the associated performance benefits.
There has also been a recent proliferation of hardware-accelerated and JAX-based RL environments. Isaac gym Makoviychuk et al. (2021) provides a GPU-accelerated simulator for a range of robotics platforms and CuLE Dalton and frosio (2020) is a CUDA reimplementation of the Atari Learning Environment Bellemare et al. (2013). Both of these environments are GPU-specific and cannot be extended to other hardware accelerators. Madrona Shacklett et al. (2023) is an extensible game-engine written in C++ that allows for GPU acceleration and parallelisation across environments. However, it requires environment code to be written in C++, limiting its accessibility. VMAS Bettini et al. (2022) provides a vectorized 2D physics engine written in PyTorch and a set of challenging multi-robot scenarios, including those from the MPE environment. For RL environments implemented in JAX, Jumanji Bonnet et al. (2023) features mostly single-agent environments with a strong focus on combinatorial problems. The authors also provide an actor-critic baseline in addition to random actions. PGX Koyamada et al. (2023) includes several board-game environments written in JAX. Gymnax Lange (2022) provides JAX implementations of the BSuite Osband et al. (2019), classic continuous control, MinAtar Young and Tian (2019) and other assorted environments. Gymnax’s sister-library, gymnax-baselines, provides PPO and ES baselines. Further extensions to Gymnax Lu et al. (2023a) also include POPGym environments Morad et al. (2023). Brax Freeman et al. (2021) reimplements the MuJoCo simulator in JAX and also provides a PPO implementation as a baseline. Jax-LOB Frey et al. (2023) implements a vectorized limit order book as an RL environment that runs on the accelerator. Perhaps the most similar to our work is Mava Pretorius et al. (2021), which provides a MAPPO baseline, as well as integration with the Robot Warehouse environment. However, none of these libraries combine a range of JAX-based MARL environments with both value-based and actor-critic baselines.
Broadly, no other work provides implementations of a wide range of hardware-accelerated MARL environments, while also implementing value-based and actor-critic baselines. Secondly, no other JAX simplification of SMAC exists. All other versions are either tied to the StarCraft II simulator or not hardware accelerated.
7. Conclusion
Hardware acceleration offers important opportunities for MARL research by lowering computational barriers, increasing the speed at which ideas can be iterated, and allowing for more thorough evaluation. We present JaxMARL, an open-source library of popular MARL environments and baseline algorithms implemented in JAX. We combine ease of use with hardware accelerator enabled efficiency to give significant speed-ups compared to traditional CPU-based implementations. Furthermore, by bringing together a wide range of MARL environments under one codebase, we have the potential to help alleviate issues with MARL’s evaluation standards. We hope that JaxMARL will help advance MARL by improving the ability of academic labs to conduct research with thorough, fast, and effective evaluations.
8. Author Contributions
This project is a large-scale effort spanning many labs and contributors.
AR11footnotemark: 122footnotemark: 2 led the design of the JaxMARL API and interface the implementation of IPPO and MPE environments. BE11footnotemark: 122footnotemark: 2 led the design and implementation of the SMAX environments and IPPO evaluations. AR and BE also led the writing of this manuscript. MG11footnotemark: 122footnotemark: 2 led the implementation of the off-policy MARL algorithms, their evaluations., and the implementation of the Switch Riddle environment.
JC11footnotemark: 1 led the implementation of the Hanabi environment and heavily assisted with benchmarking and verifying its performance. AL11footnotemark: 1 led the implementation of the Overcooked environments. GI11footnotemark: 1 led the implementation of the Multi-Agent Brax environments. TW11footnotemark: 1 led the implementation of the STORM environments. AK and AS worked on the STORM environments. CSW led the implementation of the Predator-Prey environment.
CSW, SB, MS, MJ, and RL provided invaluable discussions for project planning and implementations across the project. SB helped initiate the project plan. MS worked on the Multi-Agent Brax environments. MJ worked on the Overcooked and Hanabi environments. RL assisted with the design of the API and testing infrastructure.
SW, BL, NH, and TR provided invaluable feedback on the project, manuscript, and results.
CL11footnotemark: 122footnotemark: 2 initiated the project and led the organizational and planning efforts, speed-based benchmarking, and Coin Game implementation.
JF is the primary advisor for the project.
Appendix A Further Details on Environments
A.1. SMAX
Observations in SMAX are structured similarly to SMAC. Each agent observes the health, previous action, position, weapon cooldown and unit type of all allies and enemies in its sight range. Like SMACv2Ellis et al. (2022), we use the sight and attack ranges as prescribed by StarCraft II rather than the fixed values used in SMAC.
SMAX and SMAC have different returns. SMAC’s reward function, like SMAX’s, is split into two parts: one part for depleting enemy health, and another for winning the episode. However, in SMAC, the part which rewards depleting enemy health scales with the number of agents. This is most clearly demonstrated in 27m_vs_30m, where a random policy gets a return of around out of a maximum of because almost all the reward is for depleting enemy health or killing agents, rather than winning the episode. In SMAX, however, 50% of the total return is always for depleting enemy health, and 50% for winning. Unlike StarCraft II, where all actions happen in a randomised order in the game loop, some actions in SMAX are simultaneous, meaning draws are possible. In this case both teams get reward.
Like SMAC, each environment step in SMAX consists of eight individual time ticks. SMAX uses a discrete action space, consisting of movement in the four cardinal directions, a stop action, and a shoot action per enemy.
SMAX makes three notable simplifications of the StarCraft II dynamics to reduce complexity. First, zerg units do not regenerate health. This health regeneration is slow at health per second, and so likely has little impact on the game. Protoss units also do not have shields. Shields only recharge after 10 seconds out of combat, and therefore are unlikely to recharge during a single micromanagement task. Protoss units have additional health to compensate for their lost shields. Finally, the available unit types are reduced compared to SMAC. SMAX has no medivac, colossus or baneling units. Each of these unit types has special mechanics that were left out for the sake of simplicity. For the SMACv2 scenarios, the start positions are generated as in SMACv2, with the small difference that the ‘surrounded’ start positions now treat allies and enemies identically, rather than always spawning allies in the middle of the map. This symmetry guarantees that a 50% win rate is always achievable.
Collisions are handled by moving agents to their desired location first and then pushing them out from one another.
A.2. Coin Game
Two agents, ‘red’ and ‘blue’, move in a wrap-around grid and collect red and blue coloured coins. When an agent collects any coin, the agent receives a reward of . However, when ‘red’ collects a blue coin, ‘blue’ receives a reward of and vice versa. Once a coin is collected, a new coin of the same colour appears at a random location within the grid. If a coin is collected by both agents simultaneously, the coin is duplicated and both agents collect it. Episodes are of a set length.
A.3. Spatial-Temporal Representations of Matrix Games (STORM)
This environment features directional agents within an 8x8 grid-world with a restricted field of view. Agents cannot move backwards or share the same location. Collisions are resolved by either giving priority to the stationary agent or randomly if both are moving. Agents collect two unique resources: cooperate and defect coins. Once an agent picks up any coin, the agent’s colour shifts, indicating its readiness to interact. The agents can then release an interact beam directly ahead; when this beam intersects with another ready agent, both are rewarded based on the specific matrix game payoff matrix. The agents’ coin collections determine their strategies. For instance, if an agent has 1 cooperate coin and 3 defect coins, there’s a 25% likelihood of the agent choosing to cooperate. After an interaction, the two agents involved are frozen for five steps, revealing their coin collections to surrounding agents. After five steps, they respawn in a new location, with their coin count set back to zero. Once an episode concludes, the coin placements are shuffled. This grid-based approach to matrix games can be adapted for n-player versions. While STORM is inspired by MeltingPot 2.0, there are noteworthy differences:
Meltingpot uses pixel-based observations while we allow for direct grid access.
Meltingpot’s grid size is typically 23x15, while ours is 8x8.
Meltingpot features walls within its layout, ours does not.
Our environment introduces stochasticity by shuffling the coin placements, which remain static in Meltingpot.
Our agents begin with an empty coin inventory, making it easier for them to adopt pure cooperate or defect tactics, unlike in Meltingpot where they start with one of each coin.
MeltingPot is implemented in Lua (Ierusalimschy, 2006) where as ours is a vectorized implementation in Jax.
We deem the coin shuffling especially crucial because even large environments representing POMDPs, such as SMAC, can be solved without the need for memory if they lack sufficient randomness (Ellis et al., 2022).
A.4. Hanabi
There are a few details that differ between our Hanabi implementation and the original Hanabi Learning Environment (HLE). The most notable of these is how we choose to represent card knowledge information in the agents’ observation. In the HLE, card knowledge is observed as a colour/rank if there has been an explicit hint about a given card. As a separate feature, implicit card knowledge is represented as possible colours/ranks if there has not been an explicit hint that indicates a given card is not that colour/rank. We, on the other hand, combine implicit and explicit card knowledge, by only maintaining a representation of implicit card knowledge, which reduces to explicit card knowledge in the event an explicit hint is given about a card. This is because all possible colours/ranks are represented as 1s, whilst all ruled out colours/ranks are represented as 0s. By giving an explicit hint, all but one colour/rank are ruled out, leaving a one-hot encoding of the explicit card knowledge. We implement card knowledge this way, because knowledge updates are implemented via tensor calculus using JAX Numpy arrays of fixed shape and data type.
Appendix B Value-Based MARL Methods and Implementation details
Key features of our framework include parameter sharing, a recurrent neural network (RNN) for agents, an epsilon-greedy exploration strategy with linear decay, a uniform experience replay buffer, and the incorporation of Double Deep -Learning (DDQN) Van Hasselt et al. (2016) techniques to enhance training stability.
Unlike PyMARL, we use the Adam optimizer as the default optimization algorithm. Below is an introduction to common value-based MARL methods.
IQL (Independent -Learners) is a straightforward adaptation of Deep -Learning to multi-agent scenarios. It features multiple -Learner agents that operate independently, optimizing their individual returns. This approach follows a decentralized learning and decentralized execution pipeline.
VDN (Value Decomposition Networks) extends -Learning to multi-agent scenarios with a centralized-learning-decentralized-execution framework. Individual agents approximate their own action’s -Value, which is then summed during training to compute a jointed for the global state-action pair. Back-propagation of the global DDQN loss in respect to a global team reward optimizes the factorization of the jointed -Value.
QMIX improves upon VDN by relaxing the full factorization requirement. It ensures that a global operation on the total -Value () is equivalent to individual operations on each agent’s -Value. This is achieved using a feed-forward neural network as the mixing network, which combines agent network outputs to produce values. The global DDQN loss is computed using a single shared reward function and is back-propagated through the mixer network to the agents’ parameters. Hypernetworks generate the mixing network’s weights and biases, ensuring non-negativity using an absolute activation function. These hypernetworks are two-layered multi-layer perceptrons with ReLU non-linearity.
Appendix C Training Results
C.1. MABrax
The performance of IPPO on HalfCheeta and Walker is reported in Figure 9, with hyperparameters reported in Table 4.
C.2. MPE
Performance of -Learning baselines in all the MPE scenarios are reported in Figure 10. The upper row represents cooperative scenarios, with results for all our -learning baselines reported. The bottom row refers to competitive scenarios, and results for IQL are divided by agent types. Hyperparameters are given in Table 10
C.3. SMAX
The performance of IPPO in SMAX versus MAPPO in SMAC is shown in Figure 11 while the performance of our -learning baselines is reported in Figure 12. We do not report them together because their hyperparameters were tuned over a different number of timesteps. Hyperparameters for IPPO and the -learning methods are given in Tables 6 and 10 respectively.
Appendix D Hyperparameters
| Value | Ant | HalfCheetah | Walker |
| VF_COEF | 4.5 | 0.14 | 1.9 |
| ENT_COEF | |||
| LR | |||
| NUM_ENVS | 64 | – | – |
| NUM_STEPS | 300 | – | – |
| TOTAL_TIMESTEPS | – | – | |
| NUM_MINIBATCHES | 4 | – | – |
| GAMMA | 0.99 | – | – |
| GAE_LAMBDA | 1.0 | – | – |
| CLIP_EPS | 0.2 | – | – |
| MAX_GRAD_NORM | 0.5 | – | – |
| ACTIVATION | tanh | – | – |
| ANNEAL_LR | True | – | – |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | tanh |
| ANNEAL_LR | True |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| SCALE_CLIP_EPS | False |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | relu |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | relu |
| ANNEAL_LR | True |
| NUM_FC_LAYERS | 2 |
| LAYER_WIDTH | 512 |
| Hyperparameter | Value |
| LR | |
| NUM_ENVS | |
| NUM_STEPS | |
| TOTAL_TIMESTEPS | |
| UPDATE_EPOCHS | |
| NUM_MINIBATCHES | |
| GAMMA | |
| GAE_LAMBDA | |
| CLIP_EPS | |
| ENT_COEF | |
| VF_COEF | |
| MAX_GRAD_NORM | |
| ACTIVATION | tanh |
| ANNEAL_LR | True |
| NUM_EVALS |
| Hyperparameter | Value |
| NUM_ENVS | |
| NUM_STEPS | |
| BUFFER_SIZE | |
| BUFFER_BATCH_SIZE | |
| TOTAL_TIMESTEPS | |
| AGENT_HIDDEN_DIM | |
| AGENT_INIT_SCALE | |
| EPSILON_START | |
| EPSILON_FINISH | |
| EPSILON_ANNEAL_TIME | |
| MIXER_EMBEDDING_DIM* | |
| MIXER_HYPERNET_HIDDEN_DIM* | |
| MIXER_INIT_SCALE* | |
| MAX_GRAD_NORM | |
| TARGET_UPDATE_INTERVAL | |
| LR | |
| EPS_ADAM | |
| WEIGHT_DECAY_ADAM | |
| GAMMA | |
| NUM_TEST_EPISODES | |
| TEST_INTERVAL |
(* Parameters specific to QMix.)
| Hyperparameter | Value |
| NUM_ENVS | |
| NUM_STEPS | |
| BUFFER_SIZE | |
| BUFFER_BATCH_SIZE | |
| TOTAL_TIMESTEPS | |
| AGENT_HIDDEN_DIM | |
| AGENT_INIT_SCALE | |
| EPSILON_START | |
| EPSILON_FINISH | |
| EPSILON_ANNEAL_TIME | |
| MIXER_EMBEDDING_DIM* | |
| MIXER_HYPERNET_HIDDEN_DIM* | |
| MIXER_INIT_SCALE* | |
| MAX_GRAD_NORM | |
| TARGET_UPDATE_INTERVAL | |
| LR | |
| EPS_ADAM | |
| WEIGHT_DECAY_ADAM | |
| GAMMA | |
| NUM_TEST_EPISODES | |
| TEST_INTERVAL |
(* Parameters specific to QMix.)

### v5

@@
 JaxMARL: Multi-Agent RL Environments and Algorithms in JAX
-Abstract.
-Benchmarks play an important role in the development of machine learning algorithms.
-For example, research in reinforcement learning (RL) has been heavily influenced by available environments and benchmarks.
-However, RL environments are traditionally run on the CPU, limiting their scalability with typical academic compute.
-Recent advancements in JAX have enabled the wider use of hardware acceleration to overcome these computational hurdles, enabling massively parallel RL training pipelines and environments.
-This is particularly useful for multi-agent reinforcement learning (MARL) research.
-First of all, multiple agents must be considered at each environment step, adding computational burden, and secondly, the sample complexity is increased due to non-stationarity, decentralised partial observability, or other MARL challenges.
-In this paper, we present JaxMARL, the first open-source code base that combines ease-of-use with GPU enabled efficiency, and supports a large number of commonly used MARL environments as well as popular baseline algorithms.
-When considering wall clock time, our experiments show that per-run our JAX-based training pipeline is up to 12500x faster than existing approaches.
-This enables efficient and thorough evaluations, with the potential to alleviate the evaluation crisis of the field.
-We also introduce and benchmark SMAX, a vectorised, simplified version of the popular StarCraft Multi-Agent Challenge, which removes the need to run the StarCraft II game engine.
+Abstract
+Benchmarks are crucial in the development of machine learning algorithms, with available environments significantly influencing reinforcement learning (RL) research.
+Traditionally, RL environments run on the CPU, which limits their scalability with typical academic compute.
+However, recent advancements in JAX have enabled the wider use of hardware acceleration, enabling massively parallel RL training pipelines and environments.
+While this has been successfully applied to single-agent RL, it has not yet been widely adopted for multi-agent scenarios.
+In this paper, we present JaxMARL, the first open-source, Python-based library that combines GPU-enabled efficiency with support for a large number of commonly used MARL environments and popular baseline algorithms.
+Our experiments show that, in terms of wall clock time, our JAX-based training pipeline is around 14 times faster than existing approaches, and up to 12500x when multiple training runs are vectorized.
+This enables efficient and thorough evaluations, potentially alleviating the evaluation crisis in the field.
+We also introduce and benchmark SMAX, a JAX-based approximate reimplementation of the popular StarCraft Multi-Agent Challenge, which removes the need to run the StarCraft II game engine.
 This not only enables GPU acceleration, but also provides a more flexible MARL environment, unlocking the potential for self-play, meta-learning, and other future applications in MARL.
-We provide code at https://github.com/flairox/jaxmarl.
-Key words and phrases:
-Multi-Agent Reinforcement Learning, JAX, Benchmarks1. Introduction
-Benchmarks play a pivotal role in the development of new single and multi-agent reinforcement learning (MARL) algorithms by defining problems, enabling comparisons, and focusing efforts.
-For example, in recent years, Go and Chess drove the development of MuZero Schrittwieser et al.
-(2020) while decentralised StarCraft Micromanagement Foerster et al.
-(2017) and later the StarCraft Multi-Agent Challenge (SMAC, Samvelyan et al., 2019) resulted in the development of algorithms such as QMIX Rashid et al.
-(2020), a popular MARL technique.
-Data transfer between the CPU (where the environment is simulated) and the GPU (where the agents are evaluated) is a crucial bottleneck for simulation speed.
-Simulation speed in turn is vital for progress in reinforcement learning (RL) because RL algorithms often require a large number of environment interactions.
-This problem is even worse in MARL, where non-stationarity and decentralised partial observability greatly worsen the sample complexity Bernstein et al.
-(2002).
-Hardware acceleration and parallelisation are crucial to alleviating this, but current acceleration and parallelisation methods are typically not implemented in Python, reducing their accessibility for most machine learning researchers Shacklett et al.
-(2023); Weng et al.
-(2022).
-For example, the extremely efficient Hanabi library Hu and Foerster (2020) from Meta-AI research is implemented in C++ and has seen relatively little adoption by the community.
-However, recent advances in JAX Bradbury et al.
-(2018) have opened up new possibilities for using Python code directly with hardware accelerators, enabling the wider use of massively parallel RL training pipelines and environments.
-The JAX Bradbury et al.
-(2018) library provides composable function transformations, allowing for automatic vectorisation, device parallelisation, automatic differentiation and just-in-time (JIT) compilation with XLA Sabne (2020), for device-agnostic optimisation.
-Using JAX, both the environment rollouts and model training can happen on a hardware accelerator (such as a GPU or TPU), removing the cost of data transfer between devices and allowing for significant parallelisation.
-Recently, PureJaxRL Lu et al.
-(2022a, 2023b) has demonstrated the power of this end-to-end JAX-based approach; running both the environment and the model training on a GPU yields a 4000x speedup over a “traditional” pipeline with a GPU-trained policy but a CPU-based environment.
-These accelerations could substantially advance RL and MARL research by quickening the testing and iteration of ideas.
-Furthermore, they lower computational hurdles for in-depth MARL research, enabling researchers to utilise billions of frames and extract more performance from single GPUs.
-Alongside the current computational issues faced by MARL researchers, recent work also highlights issues with the evaluation standards and use of benchmarks in the MARL community.
+The code is available at https://github.com/flairox/jaxmarl.
+1 Introduction
+Benchmarks are crucial for developing new single and multi-agent reinforcement learning (MARL) algorithms.
+They define problems, enable comparisons, and focus research efforts.
+For example, the development of MuZero was driven by the challenges presented by Go and Chess [54].
+Similarly, decentralised StarCraft Micromanagement tasks [17] led to the creation of algorithms like QMIX [52], a popular MARL technique.
+In RL research, the runtime of simulations and algorithms is a critical factor affecting the efficiency, thoroughness, and feasibility of experiments.
+RL training pipelines often require a large number of environment interactions and long, expensive, experimental runs significantly impede research progress.
+Hardware acceleration and parallelization is an approach to address this: by running environments on hardware accelerators (e.g.
+GPUs), we can use many more environment instances in parallel than is feasible with a CPU, drastically improving runtime.
+However, such an approach typically requires significant engineering effort and often relies on non-Python codebases [56], which reduces its accessibility for many Machine Learning researchers where Python is the lingua franca.
+That said, recent releases, such as the JAX [7] library and PyTorch’s functorch module [22], have improved accessibility by enabling Python code to be parallelized and just-in-time compiled on hardware accelerators.
+This laid the foundation for PureJaxRL [38], which leveraged JAX to implement a parallelized approach, demonstrating that running both the environment and the model training on the same GPU yields a 10x speedup over a traditional pipeline with a GPU-trained policy but a CPU-based environment, and 4000x when multiple training runs are vectorized.
+This speedup enables new research directions [39, 28] and makes large-scale RL research more accessible [46].
+We introduce JaxMARL, which brings these benefits to multi-agent learning.
+To the best of our knowledge, JaxMARL is the first open-source, Python-based library which leverages JAX for GPU acceleration and supports a wide range of popular MARL environments (as shown in Figure 1) as well as algorithms.
+We show that MARL greatly benefits from this approach as under the traditional approach, of using CPU-based environments, experiments tend to be particularly slow due to the increased computational burden of training multiple agents simultaneously and the higher sample complexity arising from challenges like non-stationarity and decentralized partial observability.
+Utilizing an end-to-end JAX-based pipeline for MARL significantly accelerates these experiments, opening up new possibilities for research in this field.
+Alongside computational issues, MARL research also struggles with thorough evaluation standards [21].
 In particular, MARL papers typically only test on a few domains.
-Of the 75 recent MARL papers analysed by Gorsane et al.
-(2022), 50% used only one evaluation environment and a further 30% used only two.
-While SMAC and MPE Lowe et al.
-(2017), the two most used environments, have various tasks or maps, the lack of a standard set raises the risk of biased comparisons and incorrect conclusions.
+Of the 75 recent MARL papers analysed by [21], 50% used only one evaluation environment and a further 30% used only two.
+While SMAC [53] and MPE [36], the two most used environments, have various tasks or maps, the lack of a standard set raises the risk of biased comparisons and incorrect conclusions.
 This leads to environment overfitting and unclear progress markers.
-Instead, novel MARL methods should be tested on a wide range of domains to accurately evaluate their limits and enable better comparisons.
-The likely issue preventing this is the lack of a unified codebase and the computational burden of further evaluation.
-This paper presents JaxMARL, a Python library that for the first time brings together JAX implementations of eight common MARL environments under one API.
-We additionally provide JAX implementations for four state-of-the-art algorithms, allowing for end-to-end JAX-based training pipelines in a similar fashion to PureJaxRL.
-As outlined in Figure 1, we present a library with end-to-end hardware-accelerated training, simple Python implementations, and a broad range of MARL environments.
-By alleviating computational constraints, JaxMARL allows rapid evaluation of novel methods across a broad set of domains, and hence has the potential to be a powerful tool to address MARL’s evaluation crisis.
-Specifically, we find that JaxMARL achieves over 12500x speedup compared to “conventional” aproaches.
-We also create SMAX, a JAX-based simplification of the centralised training with decentralised execution (CTDE) benchmarks SMAC Samvelyan et al.
-(2019) and SMACv2 Ellis et al.
-(2022).
-SMAX features simplified dynamics, greater flexibility and a more sophisticated but fully-decentralised heuristic AI, while retaining the high-dimensional observation space, complex unit type interactions and procedural scenario generation that lend SMAC and SMACv2 much of their difficulty.
-As shown in Figure 2, in addition to SMAX, our library includes the most popular environments from several MARL settings.
-For centralised training with decentralised execution (CTDE), we include the Multi-Agent Particle Environments (MPE) Lowe et al.
-(2017), and Multi-Agent Brax (MABrax).
-Meanwhile, for zero-shot coordination (ZSC) and ad-hoc teamplay, we include Hanabi and Overcooked.
-Lastly, from the general-sum literature, we include the CoinGame and Spatial-Temporal Representations of Matrix Games (STORM), a representation of matrix games as grid-world scenarios with temporally extended actions.
-JaxMARL provides the first JAX implementation of these environments and unifies them in a single codebase.
-We additionally provide JAX implementations of Independent PPO (IPPO) Schulman et al.
-(2017); de Witt et al.
-(2020), QMIX, VDN Sunehag et al.
-(2017) and Independent -Learning (IQL) Mnih et al.
-(2015), four of the most common MARL algorithms, allowing new techniques to be easily benchmarked against existing practices.
-We will extend this list before the camera-ready copy, e.g.
-with the popular MAPPO Yu et al.
-(2022) algorithm.
-2. Background
-2.1. Hardware Accelerated Environments
+By alleviating computational constraints, JaxMARL allows for rapid evaluations across a broad set of environments and hence is a powerful tool to address MARL’s current evaluation crisis.
+More specifically, in this paper, our contributions are as follows:
+JAX Implementations of Popular MARL Environments: We implement a wide range of popular MARL environments in JAX, enabling fast experimentation across diverse environments.
+Taking advantage of large scale parallelism, many of our environments run over three orders of magnitude faster on a GPU than their CPU-based counterparts.
+They are implemented in Python ensuring ease-of-use, following our philosophy set out in Figure 2.
+New MARL Environment Suites: We introduce two new MARL environment suites: SMAX and STORM.
+SMAX is an approximate reimplementation of the popular SMAC(v2) [15] benchmark entirely in JAX, rather than using the StarCraft II game engine like SMAC.
+It is therefore more customizable and significantly faster: SMAX training is 40,000x faster than the equivalent SMAC implementation on a single NVIDIA 2080 when multiple training runs are vectorized.
+STORM is a general-sum environment suite inspired by the Melting Pot [34] environment suite that features temporally extended actions within social dilemmas.
+Implementation of Popular MARL Algorithms in JAX: We implement many popular MARL algorithms in JAX, such as IPPO, MAPPO and QMIX.
+As outlined in Figure 3, our training pipeline is up to 14x faster than current popular approaches, and up to 12500x when multiple training runs are vectorized.
+Comprehensive Benchmarking: We thoroughly benchmark the speed and correctness of our environments and algorithms, comparing them to existing popular repositories.
+Generally, our end-to-end JAX implementations run several thousand times faster than their CPU-based counterparts while maintaining equivalent agent performance.
+Environment Evaluation Recommendations and Best Practice: Finally, we provide environment evaluation recommendations for different MARL research settings, such as centralized training with decentralized execution and zero-shot coordination.
+We also provide scripts for large scale evaluation and plotting based on best-practice in the field.
+2 Background
+Our work brings the benefits of hardware acceleration to multi-agent reinforcement learning.
+Hardware Accelerated Environments
 JAX enables the use of Python code with any hardware accelerator, allowing researchers to write hardware-accelerated code easily.
 Within the RL community, writing environment code in JAX has gained recent popularity.
-This brings two chief advantages: firstly, environments written in JAX can be very easily parallelised by using JAX’s vmap operation, which vectorises a function across an input dimension, and secondly writing the environment in JAX allows the agent and environment to be co-located on the GPU, which eliminates the time taken to copy between CPU and GPU memory.
-Combined, these two factors bring significant increases in training speed, with PureJaxRL Lu et al.
-(2022a) achieving a x speedup over traditional training in single-agent settings.
-2.2. SMAC
+This brings three chief advantages.
+Firstly, environments written in JAX can be very easily parallelised by using JAX’s vmap operation, which vectorises a function across an input dimension.
+Secondly writing the environment in JAX allows the agent and environment to be co-located on the GPU, which eliminates the time taken to copy between CPU and GPU memory.
+Finally, the code written can be compiled just-in-time, thereby improving performance.
+Combined, these factors bring significant increases in training speed, with PureJaxRL [38] achieving a x speedup with vectorized training over traditional training in single-agent settings.
+Multi-Agent Reinforcement Learning Settings
+Multi-Agent Reinforcement Learning is a subfield of reinforcement learning that focuses on environments where multiple agents interact and learn simultaneously.
+MARL encompasses various settings that address interactions between multiple agents.
+The most widely-studied MARL setting is the fully-cooperative one, in which agents work together to achieve a common goal.
+One key framework is centralized training with decentralized execution (CTDE) [37], which allows agents to share information during the learning phase or use additional environment information.
+During execution, however, agents act based on their independent and partial observations of the environment.
+Another is zero-shot coordination, which focuses on training agents to coordinate successfully with unseen partners or in new environments without additional training [30].
+Beyond fully-cooperative settings, there are zero-sum and general-sum [34] benchmarks that study competitive and mixed incentive interactions.
+3 JaxMARL
+We present JaxMARL, a library containing simple and accessible JAX implementations of popular MARL environments and algorithms.
+JaxMARL enables significant acceleration and parallelisation over existing implementations.
+To the best of our knowledge, JaxMARL is the first open-source library that provides JAX-based implementations of a wide range of both MARL environments and baselines.
+JaxMARL’s interface is inspired by PettingZoo [61] and Gymnax [32], while the design philosophy is based on PureJaxRL [38] and CleanRL [26].
+We designed it to be a simple and easy-to-use interface for a wide range of MARL problems.
+A full specification is provided in the Appendix.
+3.1 Environments
+JaxMARL contains a diverse range of JAX reimplementations of existing environments.
+It also introduces SMAX, a novel SMAC-like JAX environment, and STORM, an expansion of matrix games to grid-world scenarios.
+In this section, we introduce our environments while further details on their implementations can be found in the Appendix.
+We measure the speed of our environments in steps per second when using random actions and compare our speed to that of the original environments in Table 3, see the Appendix for details.
+3.1.1 New Environments
+SMAX
+The StarCraft Multi-Agent Challenge (SMAC) is a popular benchmark in cooperative multi-agent reinforcement learning (MARL) but has several limitations.
+SMAC’s environment lacks sufficient stochasticity for complex policies [15], and its reliance on the StarCraft II engine makes it slow and memory-intensive [43].
+Additionally, StarCraft II’s constraints limit scenario variety and do not support competitive self-play without significant engineering.
+To address these issues, we introduce SMAX, a SMAC-like, hardware-accelerated, customizable environment.
+SMAX features more lightweight dynamics and a less exploitable AI.
+SMAX incorporates original SMAC scenarios and scenarios similar to those in SMACv2, but is also far more customizable.
+We provide more details on SMAX and how it improves on SMAC and SMACv2 in the Appendix.
+Spatial-Temporal Representations of Matrix Games (STORM)
+Inspired by the “in the Matrix” games in Melting Pot 2.0 [1], the STORM [29] environment expands on matrix games by representing them as grid-world scenarios.
+Agents collect resources which define their strategy during interactions and are rewarded based on a pre-specified payoff matrix.
+STORM can represent cooperative, competitive or general-sum games, like the prisoner’s dilemma [57].
+Thus, STORM can be used for studying paradigms such as opponent shaping, where agents act with the intent to change other agents’ learning dynamics, which has been empirically shown to lead to more prosocial outcomes [18, 65, 41, 29, 69].
+Compared to the Coin Game or simple matrix games, the grid-world setting presents a variety of new challenges such as partial observability, multi-step agent interactions, temporally-extended actions, and longer time horizons.
+Unlike the “in the Matrix” games from Melting Pot, STORM features stochasticity, increasing the difficulty [15].
+3.1.2 Existing Environments
+We have also provided JAX-based implementations of several existing environments.
+Hanabi [3] is a fully-cooperative partially-observable multiplayer card game, where players can observe others’ cards but not their own.
+It is a common benchmark for zero-shot coordination, theory of mind, and ad-hoc teamplay research [23, 24, 9, 39].
+Overcooked is commonly used for assessing fully-cooperative and fully-observable Human-AI task performance.
+Our implementation mimics the original from Overcooked-AI [10].
+For a discussion on this environment’s limitations see [33].
+MABrax is a derivative of Multi-Agent MuJoCo [49], an extension of the MuJoCo Gym environment [62] that is commonly used for benchmarking continuous multi-agent robotic control.
+Multi-Agent Particle Environment (MPE) tasks feature a 2D world with simple physics where particle agents can move, communicate, and interact with fixed landmarks [36].
+Coin Game is a two-player grid-world environment which emulates social dilemmas such as the iterated prisoner’s dilemma [57].
+While this is a common benchmark for the general-sum setting, previous work [29] has illustrated issues which STORM corrects.
+Switch Riddle [16] is a simple cooperative communication task included as a debugging tool.
+3.2 Algorithms
+In this section, we present our re-implementation of five well-known MARL baseline algorithms using JAX.
+All of our training pipelines are fully compatible with JAX’s jit and vmap functions, resulting in significant acceleration of the training processes, as outlined in Figure 3.
+It also enables parallelisation of training across many seeds and hyperparameters on a single GPU.
+We follow CleanRL’s philosophy of providing clear, single-file implementations [26] and provide a brief overview of the implemented baselines in the Appendix.
+PPO
+Q-learning
+4 Evaluation Recommendations
+Previous work [21] has found significant differences in the evaluation protocols between MARL research works.
+We identify four main research areas that would benefit from our library: cooperative centralised training with decentralised execution (CTDE) [16], zero-shot coordination [23], general-sum games, and cooperative continuous action methods.
+To aid comparisons between methods, we recommend standard minimal sets of evaluation environments for each of these settings in Table 1.
+It’s important to note that these are minimal and we encourage as broad an evaluation as possible.
+For example, in the zero-shot coordination setting, all methods should be able to evaluate on Hanabi and Overcooked.
+However, it may also be possible to evaluate such methods on the SMACv2 settings of SMAX.
+Similarly, SMAX could be used to evaluate two-player zero-sum methods by training in self-play.
+For some settings, such as continuous action environments and general-sum games, there is only one difficult environment.
+We encourage further development of JAX-based environments in these settings to improve the quality of evaluation.
+| Setting | Recommended Environments |
+| CTDE | SMAX (all scenarios), Hanabi (2-5 players), Overcooked |
+| Zero-shot Coordination | Hanabi (2 players), Overcooked (5 basic scenarios) |
+| General-Sum | STORM (iterated prisoner’s dilemma), STORM (matching pennies) |
+| Cooperative Continuous Actions | MABrax |
+To compute aggregate performance statistics, we follow the recommendations of [2] and evaluate the inter-quartile mean across the different classes of environments.
+To do this, we recommend normalising performance of the algorithms on the relevant classes of environment (for example via looking at the maximum and minimum performance across algorithms as discussed in [21]), and computing a mean per seed.
+Then compute the inter-quartile mean across aggregated statistics in each environment.
+We provide code for performing this calculation.
+This allows environment classes to be compared fairly, without over-weighting those with more individual scenarios.
+5 Results
+To demonstrate the utility of our library, we evaluate PPO against Q-Learning algorithms on a range of cooperative environments.
+We discover that not only does it have improved performance, but also that it is more practical to use for end-to-end GPU training.
+We then evaluate the speed of our library.
+We compare algorithm and environment runtimes with similar CPU-based environments.
+We find that when training PPO, JaxMARL is 31x quicker on SMAX when compared to training in SMAC and 14x quicker on MPE for a single run, and 12,500x for vectorized training runs.
+Finally, we verify the correctness of our implementations by performing thorough comparisons with prior work.
+5.1 Multi-Environment Comparison
+We provide a preliminary comparison of our PPO and Q-Learning baselines in Figure 4.
+The IQM and mean were aggregated across 9 SMAX tasks, excluding the two maps with more than 10 units, all 5 Overcooked maps, and the 2 cooperative scenarios of MPE, running 10 seeds per task.
+We did not evaluate on Hanabi or all SMAX tasks because of the large memory overhead of storing the replay buffer for the Q-Learning methods on the GPU.
+We normalize the scores of each run on each task against the highest score obtained by any algorithm in that task, and then average the scores in each environment to avoid bias towards SMAX (which contributes more tasks).111For PPO, we used MAPPO as the baseline algorithm, except for Overcooked where we used IPPO.
+For Q-Learning, we used VDN, except for SMAX where we use QMIX as it performs slightly better.
+| Environment | PPO | Q-Learning |
+| SMAX | 10 | 60 |
+| Overcooked | 1 | 10 |
+| MPE | 1.5 | 2 |
+| Hanabi | 450 | - |
+5.2 Speed Benchmarking
+We compare the performance of our environments in steps per second when using random actions to the original environments in Table 3, with details of this test provided in the Appendix.
+| Environment | Original, 1 Env | Jax, 1 Env | Jax, 100 Envs | Jax, 10k Envs |
+| MPE Simple Spread | ||||
+| Switch Riddle | ||||
+| Hanabi | ||||
+| Overcooked | ||||
+| MABrax Ant 4x2 | ||||
+| Starcraft 2s3z | ||||
+| Starcraft 27m vs 30m | ||||
+| STORM | – | |||
+| Coin Game |
+We next compare the speed of our training pipeline to that of PyMARL.
+As shown in Figure 5(a), a single Q-Learning training run for MPE’s simple spread task takes 130 seconds with JaxMARL while PyMARL requires over an hour.
+Furthermore, using JAX we can parallelise over the entire training process within a single hardware accelerator.
+For QMIX on MPE, this allows us to complete 1024 individual training runs in 198.4 seconds, compared to 1 hour and 10 minutes for a single training run with PyMARL, a speed up of 21,500x per agent.
+This analysis is repeated for IPPO in the Appendix and we find a speedup of 12,500x.
+Figure Figure 5(c) demonstrates the speedup gained from using SMAX with JaxMARL’s IPPO implementation compared to training on SMAC with PyMARL.
+Across a varying number of environment rollout threads, JaxMARL gives a speedup of up to 31x.
+5.3 Algorithm and Environment Correctness
+In this section, we compare our environment and algorithm implementations with prior work and demonstrate equivalence where applicable.
+Overcooked
+The transition dynamics of our Overcooked implementation match those of the Overcooked-AI implementation.
+We demonstrate this by training an IPPO policy on our implementation and evaluating the policy on both our implementation and the original at regular intervals.
+The performance is similar across the implementations.
+Results can be found in the Appendix.
+SMAX
+SMAX and SMAC are different environments, as they have different opponent policies and dynamics.
+However, we demonstrate some similarity between them by comparing our IPPO and MAPPO implementations against MAPPO results on SMAC, using the implementation from [58].
+We show this figure, along with a more in-depth description of their differences, in the appendix.
+We additionally present aggregate and detailed performance across SMAX in the Appendix.
+The PPO-based IPPO and MAPPO perform better than the Q-Learning methods, with the centralised information provided in the state helping MAPPO significantly outperform IPPO.
+MPE
+Our MPE environment corresponds exactly to the PettingZoo implementation.
+We validate this for each environment using a uniform-random policy on rollouts, ensuring all observations and rewards are within a tolerance of at each transition.
+We additionally compare the results of our Q-Learning and PPO implementations with existing libraries, the results of which, along with the performance of IQL on the remaining MPE environments, can be found in the Appendix.
+We compare Q-Learning and PPO on Simple Spread in Figure 6(a).
+MABrax
+As Brax differs subtly from MuJoCo, MABrax does not correspond to MAMuJoCo but the learning dynamics are qualitatively similar.
+Results therefore are not directly comparable across the two environments.
+We report mean training return across 10 seeds for IPPO on halfcheetah_6x1 in Figure 6(b).
+We additionally report the training curves for IPPO on ant_4x2, hopper_3x1, walker2d_2x3 and humanoid_9|8 in the Appendix.
+Hanabi
+Our implementation matches the Hanabi Learning Environment.
+To verify the environment’s accuracy, we obtained 10,000 action trajectories from the original C++ repository using their pretrained models.
+We confirmed that processing these action trajectories with JaxMARL produces the same states and returns as the C++ repository.
+In addition, we transferred the models trained with Pytorch/C++ to Jax and verified they obtain similar scores in JaxMARL.
+Finally, as shown in Figure 6(c), our IPPO and MAPPO models attain scores and respectively, which are better than (for IPPO) or similar to (for MAPPO) the scores attained in [67].
+6 Related Work
+MARL Libraries and Algorithms
+Several open-source libraries exist for both MARL algorithms and environments.
+The popular library PyMARL [53] provides PyTorch implementations of QMIX, VDN and IQL and integrates easily with SMAC.
+E-PyMARL [48] extends this by adding the actor-critic algorithms MADDPG [36], MAA2C [44], IA2C [44], and MAPPO, and supports SMAC, Gym [8], Robot Warehouse [11], Level-Based Foraging [11], and MPE environments.
+Recently released MARLLib [25] is instead based on the open-source RL library RLLib [35] and combines a wide range of competitive, cooperative and mixed environments with a broad set of baseline algorithms.
+Meanwhile, MALib [70] focuses on population-based MARL across a wide range of environments.
+However, none of these frameworks feature hardware-accelerated environments and thus lack the associated performance benefits.
+Hardware-Accelerated and JAX-Based RL
+There has also been a recent proliferation of hardware-accelerated and JAX-based RL environments.
+Isaac gym [42] provides a GPU-accelerated simulator for a range of robotics platforms and CuLE [12] is a CUDA reimplementation of the Atari Learning Environment [4].
+Both of these environments are GPU-specific and cannot be extended to other hardware accelerators.
+Madrona [56] is an extensible game-engine written in C++ that allows for GPU acceleration and parallelisation across environments.
+However, it requires environment code to be written in C++, limiting its accessibility.
+VMAS [5] provides a vectorized 2D physics engine written in PyTorch and a set of challenging multi-robot scenarios, including those from the MPE environment.
+For RL environments implemented in JAX, Jumanji [6] features mostly single-agent environments with a strong focus on combinatorial problems.
+The authors also provide an actor-critic baseline in addition to random actions.
+PGX [31] includes several board-game environments written in JAX.
+Gymnax [32] provides JAX implementations of the BSuite [47], classic continuous control, MinAtar [66] and other assorted environments.
+Gymnax’s sister-library, gymnax-baselines, provides PPO and ES baselines.
+Further extensions to Gymnax [40] also include POPGym environments [45].
+Brax [19] reimplements the MuJoCo simulator in JAX and also provides a PPO implementation as a baseline.
+Jax-LOB [20] implements a vectorized limit order book as an RL environment that runs on the accelerator.
+Perhaps the most similar to our work is Mava [50], which provides a MAPPO baseline, as well as integration with the Robot Warehouse environment.
+None of these libraries combine a range of JAX-based MARL environments with both value-based and actor-critic baselines.
+7 Conclusion
+Hardware acceleration offers important opportunities for MARL research by lowering computational barriers, increasing the speed at which ideas can be iterated, and allowing for more thorough evaluation.
+We present JaxMARL, an open-source library of popular MARL environments and baseline algorithms implemented in JAX.
+We combine ease of use with hardware accelerator enabled efficiency to give significant speed-ups compared to traditional CPU-based implementations.
+Furthermore, by bringing together a wide range of MARL environments under one codebase, we have the potential to help alleviate issues with MARL’s evaluation standards.
+We hope that JaxMARL will help advance MARL by enabling researchers to conduct research with thorough, fast, and effective evaluations.
+Limitations and Future Work.
+While our work provides significant advancements, several limitations remain.
+First, we observe that the speedups are less pronounced for off-policy, value-based methods.
+Additionally, there are inherent challenges with end-to-end JAX implementations, such as the difficulty in efficiently handling environments with a variable number of agents or those with massive observation sizes.
+Furthermore, our MARL environments largely re-implement or draw inspiration from existing environment suites, meaning they do not yet push the boundaries of MARL capabilities.
+Developing novel MARL environments that push the boundaries of current capabilities could provide new and challenging benchmarks for the community.
+8 Acknowledgements
+This work received funding from the EPSRC Programme Grant “From Sensing to Collaboration” (EP/V000748/1).
+MG was partially founded by the FPI-UPC Santander Scholarship FPI-UPC_93.
+JF is partially funded by the UKI grant EP/Y028481/1 (originally selected for funding by the ERC).
+JF is also supported by the JPMC Research Award and the Amazon Research Award.
+Appendix
+We structure the appendix as follows.
+Appendix A provides further background on SMAC, Appendix B describes the environments included within JaxMARL, and Appendix C sets out JaxMARL’s API.
+Appendix D explores JaxMARL’s Value-Based MARL methods, including relevant implementation details.
+Appendix E reports our speed comparisons while Appendix F details training and correctness results not included in the main text.
+Our hyperparameter values are listed in Appendix G.
+Appendix A Further Background on SMAC
 StarCraft is a popular environment for testing RL algorithms.
-It typically features features a centralised controller issuing commands to balance micromanagement, the low-level control of individual units, and macromanagement, the high level plans for economy and resource management.
-SMAC Samvelyan et al.
-(2019), instead, focuses on decentralised unit micromanagement across a range of scenarios divided into three broad categories: symmetric, where each side has the same units, asymmetric, where the enemy team has more units, and micro-trick, which are scenarios designed specifically to feature a particular StarCraft micromanagement strategy.
-SMACv2 Ellis et al.
-(2022) demonstrates that open-loop policies can be effective on SMAC and adds additional randomly generated scenarios to rectify SMAC’s lack of stochasticity.
+It typically features a centralised controller issuing commands to balance micromanagement, the low-level control of individual units, and macromanagement, the high level plans for economy and resource management.
+SMAC [53], instead, focuses on decentralised unit micromanagement across a range of scenarios divided into three broad categories: symmetric, where each side has the same units, asymmetric, where the enemy team has more units, and micro-trick, which are scenarios designed specifically to feature a particular StarCraft micromanagement strategy.
+SMACv2 [15] demonstrates that open-loop policies can be effective on SMAC and adds additional randomly generated scenarios to rectify SMAC’s lack of stochasticity.
 However, both of these environments rely on running the full game of StarCraft II, which severely increases their CPU and memory requirements.
-SMAClite Michalski et al.
-(2023) attempts to alleviate this computational burden by recreating the SMAC environment primarily in NumPy, with some core components written in C++.
+SMAClite [43] attempts to alleviate this computational burden by recreating the SMAC environment primarily in NumPy, with some core components written in C++.
 While this is much more lightweight than SMAC, it cannot be run on a GPU and therefore cannot be parallelised effectively with typical academic hardware, which commonly has very few CPU cores compared to industry clusters.
-3. JaxMARL
-We present JaxMARL, a library containing simple and accessible JAX implementations of popular MARL environments and algorithms.
-JAX enables significant acceleration and parallelisation over existing implementations.
-To the best of our knowledge, JaxMARLis the first open source library that provides JAX-based implementations of a wide range of MARL environments and baselines.
-3.1. API
-The interface of JaxMARL is inspired by PettingZoo Terry et al.
-(2021) and Gymnax.
+Appendix B Further Details on Environments
+B.1 SMAX
+The StarCraft Multi-Agent Challenge (SMAC) is a popular benchmark but has a number of shortcomings.
+First, as noted and addressed in SMACv2, SMAC is not particularly stochastic.
+This means that non-trivial win-rates are possible on many SMAC maps by conditioning a policy only on the timestep and agent ID.
+Additionally, SMAC relies on StarCraft II as a simulator.
+While this allows SMAC to use the wide range of units, objects and terrain available in StarCraft II, running an entire instance of StarCraft II is slow and memory intensive.
+StarCraft II runs on the CPU and therefore SMAC’s parallelisation is severely limited with typical academic compute.
+Using the StarCraft II game engine also constrains environment design.
+For example, StarCraft II groups units into three races and does not allow units of different races on the same team, limiting the variety of scenarios that can be generated.
+Secondly, SMAC does not support a competitive self-play setting without significant engineering work.
+The purpose of SMAX is to address these limitations.
+It provides access to a simplified SMAC-like, hardware-accelerated, customisable environment that supports self-play and custom unit types.
+SMAX models units as discs in a continuous 2D space.
+As listed in Table 4, we include all SMAC(v1) scenarios alongside three inspired by SMAC(v2).
+Observations in SMAX are structured similarly to SMAC.
+Each agent observes the health, previous action, position, weapon cooldown and unit type of all allies and enemies in its sight range.
+Like SMACv2[15], we use the sight and attack ranges as prescribed by StarCraft II rather than the fixed values used in SMAC.
+SMAX and SMAC have different returns.
+SMAC’s reward function, like SMAX’s, is split into two parts: one part for depleting enemy health, and another for winning the episode.
+However, in SMAC, the part which rewards depleting enemy health scales with the number of agents.
+This is most clearly demonstrated in 27m_vs_30m, where a random policy gets a return of around out of a maximum of because almost all the reward is for depleting enemy health or killing agents, rather than winning the episode.
+In SMAX, however, 50% of the total return is always for depleting enemy health, and 50% for winning.
+SMAX also features a different, and more sophisticated, heuristic AI.
+The heuristic in SMAC simply moves to a fixed location, attacking any enemies it encounters along the way, and the heuristic in SMACv2 globally pursues the nearest agent.
+Thus the SMAC AI often does not aggressively pursue enemies that run away, and cannot generalise to the SMACv2 start positions, whereas the SMACv2 heuristic AI conditions on global information and is exploitable because of its tendency to flip-flop between two similarly close enemies.
+SMAC’s heuristic AI must be coded in the map editor, which does not provide a simple coding interface.
+Figure 7 demonstrates these limitations.
+In contrast, SMAX features a decentralised heuristic AI that can effectively find enemies without requiring the global information of the SMACv2 heuristic.
+This guarantees that in principle a 50% win rate is always achievable by copying the decentralised heuristic policy exactly.
+This means any win-rate below 50% represents a concrete failure to learn.
+Some of the capabilities of the SMAX heuristic AI are illustrated in the Figure below.
+Unlike StarCraft II, where all actions happen in a randomised order in the game loop, some actions in SMAX are simultaneous, meaning draws are possible. In this case both teams get reward.
+Like SMAC, each environment step in SMAX consists of eight individual time ticks.
+SMAX uses a discrete action space, consisting of movement in the four cardinal directions, a stop action, and a shoot action per enemy.
+SMAX makes three notable simplifications of the StarCraft II dynamics to reduce complexity.
+First, zerg units do not regenerate health.
+This health regeneration is slow at health per second, and so likely has little impact on the game.
+Protoss units also do not have shields.
+Shields only recharge after 10 seconds out of combat, and therefore are unlikely to recharge during a single micromanagement task.
+Protoss units have additional health to compensate for their lost shields.
+Finally, the available unit types are reduced compared to SMAC.
+SMAX has no medivac, colossus or baneling units.
+Each of these unit types has special mechanics that were left out for the sake of simplicity.
+For the SMACv2 scenarios, the start positions are generated as in SMACv2, with the small difference that the ‘surrounded’ start positions now treat allies and enemies identically, rather than always spawning allies in the middle of the map.
+This symmetry guarantees that a 50% win rate is always achievable.
+Collisions are handled by moving agents to their desired location first and then pushing them out from one another.
+| Scenario | Ally Units | Enemy Units | Start Positions |
+| 2s3z | 2 stalkers and 3 zealots | 2 stalkers and 3 zealots | Fixed |
+| 3s5z | 3 stalkers and 5 zealots | 3 stalkers and 5 zealots | Fixed |
+| 5m_vs_6m | 5 marines | 6 marines | Fixed |
+| 10m_vs_11m | 10 marines | 11 marines | Fixed |
+| 27m_vs_30m | 27 marines | 30 marines | Fixed |
+| 3s5z_vs_3s6z | 3 stalkers and 5 zealots | 3 stalkers and 6 zealots | Fixed |
+| 3s_vs_5z | 3 stalkers | 5 zealots | Fixed |
+| 6h_vs_8z | 6 hydralisks | 8 zealots | Fixed |
+| smacv2_5_units | 5 uniformly randomly chosen | 5 uniformly randomly chosen | SMACv2-style |
+| smacv2_10_units | 10 uniformly randomly chosen | 10 uniformly randomly chosen | SMACv2-style |
+| smacv2_20_units | 20 uniformly randomly chosen | 20 uniformly randomly chosen | SMACv2-style |
+B.2 Spatial-Temporal Representations of Matrix Games (STORM)
+This environment features directional agents within an 8x8 grid world with a restricted field of view.
+For a visual description, see Figure 9.
+Agents cannot move backwards or share the same location.
+Collisions are resolved by either giving priority to the stationary agent or randomly if both are moving.
+Agents collect two unique resources: cooperate and defect coins.
+Once an agent picks up any coin, the agent’s colour shifts, indicating its readiness to interact.
+The agents can then release an interact beam directly ahead; when this beam intersects with another ready agent, both are rewarded based on the specific matrix game payoff matrix.
+The agents’ coin collections determine their strategies.
+For instance, if an agent has 1 cooperate coin and 3 defect coins, there is a 25% likelihood of the agent choosing to cooperate.
+After an interaction, the two agents involved are frozen for five steps, revealing their coin collections to surrounding agents.
+After five steps, they respawn in a new location, with their coin count set back to zero.
+Once an episode concludes, the coin placements are shuffled.
+This grid-based approach to matrix games can be adapted for n-player versions.
+While STORM is inspired by MeltingPot 2.0, there are noteworthy differences:
+Meltingpot uses pixel-based observations while we allow for direct grid access.
+Meltingpot’s grid size is typically 23x15, while ours is 8x8.
+Meltingpot features walls within its layout, ours does not.
+Our environment introduces stochasticity by shuffling the coin placements, which remain static in Meltingpot.
+Our agents begin with an empty coin inventory, making it easier for them to adopt pure cooperate or defect tactics, unlike in Meltingpot where they start with one of each coin.
+MeltingPot is implemented in Lua [27] where as ours is a vectorized implementation in JAX.
+We deem the coin shuffling especially crucial because even large environments representing POMDPs, such as SMAC, can be solved without the need for memory if they lack sufficient randomness [15].
+B.3 Coin Game
+Coin Game is a two-player grid-world environment which emulates social dilemmas such as the iterated prisoner’s dilemma [57].
+Used as a benchmark for the general-sum setting, it expands on simpler social dilemmas by adding a high-dimensional state.
+Two players, ‘red’ and ‘blue’ move in a grid world and are each awarded 1 point for collecting any coin.
+However, ‘red’ loses 2 points if ‘blue’ collects a red coin and vice versa.
+Thus, if both agents ignore colour when collecting coins their expected reward is 0.
+Two agents, ‘red’ and ‘blue’, move in a wrap-around grid and collect red and blue coloured coins.
+When an agent collects any coin, the agent receives a reward of .
+However, when ‘red’ collects a blue coin, ‘blue’ receives a reward of and vice versa.
+Once a coin is collected, a new coin of the same colour appears at a random location within the grid.
+If a coin is collected by both agents simultaneously, the coin is duplicated and both agents collect it.
+Episodes are of a set length.
+B.4 Switch Riddle
+Originally used to illustrate the Differentiable Inter-Agent Learning algorithm [16], Switch Riddle is a simple cooperative communication environment that we include as a debugging tool.
+prisoners held by a warden can secure their release by collectively ensuring that each has passed through a room with a light bulb and a switch.
+Each day, a prisoner is chosen at random to enter this room.
+They have three choices: do nothing, signal to the next prisoner by toggling the light, or inform the warden they think all prisoners have been in the room.
+The game ends when a prisoner informs the warden or the maximum time steps are reached.
+The rewards are +1 if the prisoner informs the warden, and all prisoners have been in the room, -1 if the prisoner informs the warden before all prisoners have taken their turn, and 0 otherwise, including when the maximum time steps are reached.
+We benchmark using the implementation from [68].
+B.5 Hanabi
+Hanabi is a fully cooperative partially observable multiplayer card game, where players can observe other players’ cards but not their own.
+To win, the team must play a series of cards in a specific order while sharing only a limited amount of information between players.
+As reasoning about the beliefs and intentions of other agents is central to performance, it is a common benchmark for ZSC and ad-hoc teamplay research.
+Our implementation is inspired by the Hanabi Learning Environment [3] and includes custom configurations for varying game settings, such as the number of colours/ranks, number of players, and number of hint tokens.
+Compared to the Hanabi Learning Environment, which is written in C++ and split over dozens of files, our implementation is a single easy-to-read Python file, which simplifies interfacing with the library and running experiments.
+B.6 Overcooked
+Inspired by the popular videogame of the same name, Overcooked is commonly used for assessing fully cooperative and fully observable Human-AI task performance.
+The aim is to quickly prepare and deliver soup, which involves putting three onions in a pot, cooking the soup, and serving it into bowls.
+Two agents, or cooks, must coordinate to effectively divide the tasks to maximise their common reward signal.
+Our implementation mimics the original from Overcooked-AI [10], including all five original layouts and a simple method for creating additional ones.
+For a discussion on the limitations of the Overcooked-AI environment, see [33].
+B.7 Multi-Agent Particle Environments (MPE)
+The multi-agent particle environments feature a 2D world with simple physics where particle agents can move, communicate, and interact with fixed landmarks.
+Each specific environment varies the format of the world and the agents’ abilities, creating a diverse set of tasks that include both competitive and cooperative settings.
+We implement all the MPE scenarios featured in the PettingZoo library and the transitions of our implementation map exactly to theirs.
+We additionally include a fully cooperative predator-prey variant of simple tag, presented in [49].
+The code is structured to allow for straightforward extensions, enabling further tasks to be added.
+B.8 Multi-Agent Brax (MABrax)
+MABrax is a derivative of Multi-Agent MuJoCo [49], an extension of the MuJoCo Gym environment [62] that is commonly used for benchmarking continuous multi-agent robotic control.
+Our implementation utilises Brax[19] as the underlying physics engine and includes five of Multi-Agent MuJoCo’s multi-agent factorisation tasks, where each agent controls a subset of the joints and only observes the local state.
+The included tasks, illustrated in Figure 1, are: ant_4x2, halfcheetah_6x1, hopper_3x1, humanoid_9|8, and walker2d_2x3.
+The task descriptions mirror those from Gymnasium-Robotics [13].
+Appendix C JaxMARL’s API
+The interface of JaxMARL is inspired by PettingZoo [61] and Gymnax.
 We designed it to be a simple and easy-to-use interface for a wide-range of MARL problems.
@@
 Whenever JaxMARL environments correspond to existing CPU-based implementations, the version numbers match.
-3.2. Environments
-JaxMARL contains a diverse range of environments, all implemented in JAX.
-We also introduce SMAX, a SMAC-like environment implemented entirely in JAX.
-In this section we introduce these environments and provide details on their implementations.
-SMAX
-The StarCraft Multi-Agent Challenge (SMAC) is a popular benchmark but has a number of shortcomings.
-First, as noted and addressed in prior work Ellis et al.
-(2022), it is not sufficiently stochastic to require complex closed-loop policies.
-Additionally, SMAC relies on StarCraft II as a simulator.
-While this allows SMAC to use the wide range of units, objects and terrain available in StarCraft II, running an entire instance of StarCraft II is slow Michalski et al.
-(2023) and memory intensive.
-StarCraft II runs on the CPU and therefore SMAC’s parallelisation is severely limited with typical academic compute.
-| Scenario | Ally Units | Enemy Units | Start Positions |
-| 2s3z | 2 stalkers and 3 zealots | 2 stalkers and 3 zealots | Fixed |
-| 3s5z | 3 stalkers and 5 zealots | 3 stalkers and 5 zealots | Fixed |
-| 5m_vs_6m | 5 marines | 6 marines | Fixed |
-| 10m_vs_11m | 10 marines | 11 marines | Fixed |
-| 27m_vs_30m | 27 marines | 30 marines | Fixed |
-| 3s5z_vs_3s6z | 3 stalkers and 5 zealots | 3 stalkers and 6 zealots | Fixed |
-| 3s_vs_5z | 3 stalkers | 5 zealots | Fixed |
-| 6h_vs_8z | 6 hydralisks | 8 zealots | Fixed |
-| smacv2_5_units | 5 uniformly randomly chosen | 5 uniformly randomly chosen | SMACv2-style |
-| smacv2_10_units | 10 uniformly randomly chosen | 10 uniformly randomly chosen | SMACv2-style |
-| smacv2_20_units | 20 uniformly randomly chosen | 20 uniformly randomly chosen | SMACv2-style |
-Using the StarCraft II game engine constrains environment design.
-For example, StarCraft II groups units into three races and does not allow units of different races on the same team, limiting the variety of scenarios that can be generated.
-Secondly, SMAC does not support a competitive self-play setting without significant engineering work.
-The purpose of SMAX is to address these limitations.
-It provides access to a SMAC-like, hardware-accelerated, customisable environment that supports self-play and custom unit types.
-Units in SMAX are modelled as circles in a two-dimensional continuous space.
-SMAX makes a number of additional simplifications to the dynamics of StarCraft II, details of which are given in Appendix A.1.
-SMAX also features a different, and more sophisticated, heuristic AI.
-The heuristic in SMAC simply moves to a fixed location Michalski et al.
-(2023), attacking any enemies it encounters along the way, and the heuristic in SMACv2 globally pursues the nearest agent.
-Thus the SMAC AI often does not aggressively pursue enemies that run away, and cannot generalise to the SMACv2 start positions, whereas the SMACv2 heuristic AI conditions on global information and is exploitable because of its tendency to flip-flop between two similarly close enemies.
-SMAC’s heuristic AI must be coded in the map editor, which does not provide a simple coding interface.
-In contrast, SMAX features a decentralised heuristic AI that can effectively find enemies without requiring the global information of the SMACv2 heuristic.
-This guarantees that in principle a 50% win rate is always achievable by copying the decentralised heuristic policy exactly.
-This means any win-rate below 50% represents a concrete failure to learn.
-SMAX scenarios incorporate both a number of the original scenarios from SMAC and scenarios similar to those found in SMACv2.
-The latter sample units uniformly across all SMAX unit types (stalker, zealot, hydralisk, zergling, marine, marauder) and ensure fairness by having identical team composition for the enemy and ally teams.
-We provide more details on SMAX in Appendix A.1.
-Overcooked
-Inspired by the popular videogame of the same name, Overcooked is commonly used for assessing fully cooperative and fully observable Human-AI task performance.
-The aim is to quickly prepare and deliver soup, which involves putting three onions in a pot, cooking the soup, and serving it into bowls.
-Two agents, or cooks, must coordinate to effectively divide the tasks to maximise their common reward signal.
-Our implementation mimics the original from Overcooked-AI (Carroll et al., 2019), including all five original layouts and a simple method for creating additional ones.
-For a discussion on the limitations of the Overcooked-AI environment, see Lauffer et al.
-(2023).
-Hanabi
-Hanabi is a fully cooperative partially observable multiplayer card game, where players can observe other players’ cards but not their own.
-To win, the team must play a series of cards in a specific order while sharing only a limited amount of information between players.
-As reasoning about the beliefs and intentions of other agents is central to performance, it is a common benchmark for ZSC and ad-hoc teamplay research.
-Our implementation is inspired by the Hanabi Learning Environment Bard et al.
-(2020) and includes custom configurations for varying game settings, such as the number of colours/ranks, number of players, and number of hint tokens.
-Compared to the Hanabi Learning Environment, which is written in C++ and split over dozens of files, our implementation is a single easy-to-read Python file, which simplifies interfacing with the library and running experiments.
-Multi-Agent Particle Environments (MPE)
-The multi-agent particle environments feature a 2D world with simple physics where particle agents can move, communicate, and interact with fixed landmarks.
-Each specific environment varies the format of the world and the agents’ abilities, creating a diverse set of tasks that include both competitive and cooperative settings.
-We implement all the MPE scenarios featured in the PettingZoo library and the transitions of our implementation map exactly to theirs.
-We additionally include a fully cooperative predator-prey variant of simple tag, presented in Peng et al.
-(2021).
-The code is structured to allow for straightforward extensions, enabling further tasks to be added.
-Multi-Agent Brax (MABrax)
-MABrax is a derivative of Multi-Agent MuJoCo Peng et al.
-(2021), an extension of the MuJoCo Gym environment Todorov et al.
-(2012) that is commonly used for benchmarking continuous multi-agent robotic control.
-Our implementation utilises BraxFreeman et al.
-(2021) as the underlying physics engine and includes five of Multi-Agent MuJoCo’s multi-agent factorisation tasks, where each agent controls a subset of the joints and only observes the local state.
-The included tasks, illustrated in Figure 2, are: ant_4x2, halfcheetah_6x1, hopper_3x1, humanoid_9|8, and walker2d_2x3.
-The task descriptions mirror those from Gymnasium-Robotics de Lazcano et al. (2023).
-| Environment | Original, 1 Env | Jax, 1 Env | Jax, 100 Envs | Jax, 10k Envs | Maximum Speedup |
-| MPE Simple Spread | |||||
-| MPE Simple Reference | |||||
-| Switch Riddle | |||||
-| Hanabi | |||||
-| Overcooked | |||||
-| MABrax Ant 4x2 | |||||
-| Starcraft 2s3z | |||||
-| Starcraft 27m vs 30m | |||||
-| STORM | – | – | |||
-| Coin Game |
-Coin Game
-Coin Game is a two-player grid-world environment which emulates social dilemmas such as the iterated prisoner’s dilemma Snyder (1971).
-Used as a benchmark for the general-sum setting, it expands on simpler social dilemmas by adding a high-dimensional state.
-Two players, ‘red’ and ‘blue’ move in a grid world and are each awarded 1 point for collecting any coin.
-However, ‘red’ loses 2 points if ‘blue’ collects a red coin and vice versa.
-Thus, if both agents ignore colour when collecting coins their expected reward is 0.
-Further details are provided in Appendix A.2.
-Spatial-Temporal Representations of Matrix Games (STORM)
-Inspired by the “in the Matrix” games in Melting Pot 2.0 Agapiou et al.
-(2022), the STORM (Khan et al., 2022) environment expands on matrix games by representing them as grid-world scenarios.
-Agents collect resources which define their strategy during interactions and are rewarded based on a pre-specified payoff matrix.
-This allows for the embedding of fully cooperative, competitive or general-sum games, such as the prisoner’s dilemma Snyder (1971).
-Thus, STORM can be used for studying paradigms such as opponent shaping, where agents act with the intent to change other agents’ learning dynamics, which has been empirically shown to lead to more prosocial outcomes (Foerster et al., 2018; Khan et al., 2022; Lu et al., 2022b; Zhao et al., 2022).
-Compared to the Coin Game or matrix games, the grid-world setting presents a variety of new challenges such as partial observability, multi-step agent interactions, temporally-extended actions, and longer time horizons.
-Unlike the “in the Matrix” games from Melting Pot, STORM features stochasticity, increasing the difficulty Ellis et al.
-(2022).
-A further environment specification is provided in Appendix A.3.
-Switch Riddle
-Originally used to illustrate the Differentiable Inter-Agent Learning algorithm Foerster et al.
-(2016), Switch Riddle is a simple cooperative communication environment that we include as a debugging tool.
-prisoners held by a warden can secure their release by collectively ensuring that each has passed through a room with a light bulb and a switch.
-Each day, a prisoner is chosen at random to enter this room.
-They have three choices: do nothing, signal to the next prisoner by toggling the light, or inform the warden they think all prisoners have been in the room.
-The game ends when a prisoner informs the warden or the maximum time steps are reached.
-The rewards are +1 if the prisoner informs the warden, and all prisoners have been in the room, -1 if the prisoner informs the warden before all prisoners have taken their turn, and 0 otherwise, including when the maximum time steps are reached.
-We benchmark using the implementation from Zhang et al.
-(2022).
-3.3. Algorithms
-In this section, we present our re-implementation of four well known MARL baseline algorithms using JAX.
-The primary objective of these baselines is to provide a structured framework for developing MARL algorithms leveraging the advantages of the JaxMARL environments.
-All of the training pipelines are fully compatible with JAX’s JIT and VMAP functions, resulting in a significant acceleration of both the training and metric evaluation processes.
-This enables parallelisation of training across various seeds and hyperparameters on a single machine in parallel.
-We follow the CleanRL philosophy of providing clear, single-file implementations Huang et al.
-(2022).
-IPPO
--learning Methods
-Our -Learning baselines, including Independent -Learning (IQL) Tampuu et al.
-(2017), Value Decomposition Networks (VDN) Sunehag et al.
-(2018), and QMIX Rashid et al.
-(2018), have been implemented in accordance with the PyMARL codebase Rashid et al.
-(2018) to ensure consistency with published results and enable direct comparisons with PyTorch.
-Our baselines natively support aggregating trajectories from batched environments, simplifying parallelisation.
-This approach is more convenient than managing environments on distinct threads and subsequently aggregating results, as done in PyMARL.
-We provide a brief overview of the implemented baselines in the Appendix.
-4. Results
-In our results, we aim to demonstrate the speed and correctness of our environments and algorithms.In several cases, minor changes to the environments mean that our environments do not exactly match the originals on a step-by-step level.
-We therefore demonstrate the correctness in different ways for each environment and discuss each separately.
-By combining this evidence, we demonstrate that our library provides overall correct and far quicker baselines on a wide range of sufficiently correct and easily-modifiable environments.
-4.1. Environment Speed
-We measure the performance of our environments in steps per second when using random actions and compare to the original environments in Table 2 and Figure 4.
-All results were collected on a single NVIDIA A100 GPU and AMD EPYC 7763 64-core processor.
-Environments were rolled out for 1000 sequential steps.
-Many environments have comparable performance to JaxMARL when comparing single environments, but the ease of parallelisation with Jax allows for more efficient scaling compared to CPU-based environments.
-For example, MPE Simple Spread’s JAX implementation is 2̃0x slower than the original when comparing a single environment, but even when only running environments in parallel, the JAX environment is already over x faster.
-When considering environments, the JAX versions are much faster, achieving speedups of up to x over the single-threaded environment (in the case of Overcooked).
-Running this many environments in parallel using CPU environments would require a large CPU cluster and sophisticated communication mechanisms.
-This engineering is typically beyond the resources of academic labs, and therefore JaxMARL can unlock new research directions for such institutions.
-4.2. Algorithm Speed
-We investigate the speed of our IPPO implementation in Figure 5.
-By vectorising over agents, it is possible to train a vast number of agents in a fraction of the time it takes to train a single agent without hardware acceleration.
-For MPE, it is possible to train 1024 teams in seconds, which is less than seconds per teams of agents.
-A single run of MARLLIB’s IPPO implementation on the same hardware takes around seconds on average.
-This represents a speedup of over x.
-Our JAX-based -learning algorithms also offer significant speed advantages.
-In Figure 5(a), training a single IQL, VDN, or QMIX policy in MPE takes seconds while using PyMarl takes over an hour.
-Training 1024 QMIX learners in a batch requires seconds, which translates to seconds per learner, indicating a x speedup.
-This speedup is not as large as for IPPO because -learning baselines are typically trained with fewer parallel environments.
-In our experiments, we used 8 parallel environments for -learning compared to the 25 or 64 used for PPO.
-This difference is due to -learners benefiting more from a buffer with trajectories collected by different policies, resulting in a more frequent policy update, rather than collecting many trajectories with the same policy in parallel.
-For SMAX, we compare our vectorised IPPO baseline to the MAPPO implementation provided in Sun et al.
-(2023).
-MAPPO utilises an RNN and IPPO uses a feed forward network.
-This was run on a machine with a 64-core CPU and NVIDIA 2080Ti GPU.
-Additionally, as discussed in Section 3.2, SMAC and SMAX are different environments.
-These caveats aside, the differences in performance are so striking that we believe this clearly demonstrates the advantages of our approach.
-We trained 512 SMAX teams on 2s3z in under 33 minutes, whereas a single training run of PyTorch IPPO implementation takes 44 hours on average.
-This is roughly a x speedup.
-4.3. Algorithm Correctness
-We verify the correctness of our algorithm implementations by comparing to baselines from other libraries on the MPE Simple Spread and Simple Speaker Listener environments.
-For IPPO we report the mean return across seeds in Figure 4(b).
-Results were collected on the same hardware as listed in Section 4.1.
-Our IPPO implementation obtains the same performance as MARLLIB and runs x quicker, taking only ten seconds to train.
-For the -learning algorithms, we verify the correctness by comparing with PyMARL implementations of the same algorithms on the MPE Simple Spread and Simple Speaker Listener environments.
-IQL, VDN and QMIX all obtain the same or better results than their PyMARL counterparts.
-The returns are from greedy policies and averaged across 8 runs.
-The hyperparameters used are from the PyMARL library.
-4.4. Environment Correctness
-MPE
-Our MPE environment corresponds exactly to the PettingZoo implementation.
-We validate this for each environment using a uniform-random policy on rollouts, ensuring all observations and rewards are within a tolerance of at each transition.
-This tolerance accounts for non-determinism due to running floating point computation on the GPU.
-The correspondence is also shown through the performance of IPPO in Figure 4(b) and the -learning algorithms in Figures 5(b) and 5(c) respectively, as the performance of these algorithms is inline with existing baselines Yu et al.
-(2022).
-We additionally report training performance for IQL on the remaining MPE environments in Appendix C.2.
-Overcooked
-The transition dynamics of our Overcooked implementation match those of the Overcooked-AI implementation.
-We demonstrate this by training an IPPO policy on our implementation and evaluating the policy on both our Overcooked implementation and the original at regular intervals.
-Results are illustrated in Figure 6(a) and performance is similar, demonstrating their equivalence.
-SMAX
-SMAX and SMAC are different environments.
-However, we demonstrate some similarity between them by comparing our IPPO and MAPPO implementations against MAPPO results on SMAC, using the implementation from Sun et al.
-(2023).
-We show this in Figure 8.
-SMAX and SMAC have different opponent policies and dynamics, which makes this comparison more qualitative than precise.
-We describe the differences between the two in more depth in in the supplementary material.
-However, despite these differences, the environments seem similarly difficult, with some environments being more difficult in SMAC, and some more difficult in SMAX.
-This is shown in Figure 8 and in the supplementary material.
-MABrax
-As Brax differs subtly from MuJoCo, MABrax does not correspond to MAMuJoCo but the learning dynamics are qualitatively similar.
-To demonstrate this, we report mean training return across 10 seeds for IPPO on ant_4x2 in Figure 6(b), and our results are in line with the performance of TRPO reported in Kuba et al. (2021).
-We report the performance of IPPO on HalfCheetah and Walker in Appendix C.1, the results are also in line with TRPO.
-| Setting | Recommended Environments |
-| CTDE | SMAX (all scenarios), Hanabi (2-5 players), Overcooked |
-| Zero-shot Coordination | Hanabi (2 players), Overcooked (5 basic scenarios) |
-| General-Sum | STORM (iterated prisoner’s dilemma), STORM (matching pennies) |
-| Cooperative Continuous Actions | MABrax |
-Hanabi
-Our implementation does not correspond exactly to the Hanabi Learning Environment as we use a subtly different observation space, with the reasoning given in Appendix A.4.
-To demonstrate qualitative similarity, we train IPPO on Hanabi in self-play with 2 players, with the mean test return across 3 seeds reported in Figure 6(c).
-STORM, Coin Game & Switch Riddle
-STORM differs from Melting Pot 2.0 significantly, making direct comparisons challenging, with differences discussed in Appendix A.3.
-Furthermore, STORM and Coin Game are general-sum games, so the environment returns of IPPO in self-play would not be a good indicator of performance.
-Switch Riddle is a simple diagnostic environment – we do not use it for thorough evaluations.
-5. Evaluation Recommendations
-Previous work Gorsane et al.
-(2022) has found significant differences in the evaluation protocols between MARL research works.
-We identify four main research areas that would benefit from our library: cooperative centralised training with decentralised execution (CTDE) Foerster et al.
-(2016), zero-shot coordination Hu et al.
-(2020), general-sum games, and cooperative continuous action methods.
-To aid comparisons between methods, we recommend standard minimal sets of evaluation environments for each of these settings in Table 3.
-It’s important to note that these are minimal and we encourage as broad an evaluation as possible.
-For example, in the zero-shot coordination setting, all methods should be able to evaluate on Hanabi and Overcooked.
-However, it may also be possible to evaluate such methods on the SMACv2 settings of SMAX.
-Similarly, SMAX could be used to evaluate two-player zero-sum methods by training in self-play.
-For some settings, such as continuous action environments and general-sum games, there is only one difficult environment.
-We encourage further development of JAX-based environments in these settings to improve the quality of evaluation.
-6. Related Work
-Several open-source libraries exist for both MARL algorithms and environments.
-The popular library PyMARL Samvelyan et al.
-(2019) provides PyTorch implementations of QMIX, VDN and IQL and integrates easily with SMAC.
-E-PyMARL Papoudakis et al.
-(2021) extends this by adding the actor-critic algorithms MADDPG Lowe et al.
-(2017), MAA2C Mnih et al.
-(2016), IA2C Mnih et al.
-(2016), and MAPPO, and supports the SMAC, Gym Brockman et al.
-(2016), Robot Warehouse Christianos et al.
-(2020), Level-Based Foraging Christianos et al.
-(2020), and MPE environments.
-Recently released MARLLib Hu et al.
-(2022) is instead based on the open-source RL library RLLib Liang et al.
-(2018) and combines a wide range of competitive, cooperative and mixed environments with a broad set of baseline algorithms.
-Meanwhile, MALib Zhou et al.
-(2023) focuses on population-based MARL across a wide range of environments.
-However, none of these frameworks feature hardware-accelerated environments and thus lack the associated performance benefits.
-There has also been a recent proliferation of hardware-accelerated and JAX-based RL environments.
-Isaac gym Makoviychuk et al.
-(2021) provides a GPU-accelerated simulator for a range of robotics platforms and CuLE Dalton and frosio (2020) is a CUDA reimplementation of the Atari Learning Environment Bellemare et al.
-(2013).
-Both of these environments are GPU-specific and cannot be extended to other hardware accelerators.
-Madrona Shacklett et al.
-(2023) is an extensible game-engine written in C++ that allows for GPU acceleration and parallelisation across environments.
-However, it requires environment code to be written in C++, limiting its accessibility.
-VMAS Bettini et al.
-(2022) provides a vectorized 2D physics engine written in PyTorch and a set of challenging multi-robot scenarios, including those from the MPE environment.
-For RL environments implemented in JAX, Jumanji Bonnet et al.
-(2023) features mostly single-agent environments with a strong focus on combinatorial problems.
-The authors also provide an actor-critic baseline in addition to random actions.
-PGX Koyamada et al.
-(2023) includes several board-game environments written in JAX.
-Gymnax Lange (2022) provides JAX implementations of the BSuite Osband et al.
-(2019), classic continuous control, MinAtar Young and Tian (2019) and other assorted environments.
-Gymnax’s sister-library, gymnax-baselines, provides PPO and ES baselines.
-Further extensions to Gymnax Lu et al.
-(2023a) also include POPGym environments Morad et al.
-(2023).
-Brax Freeman et al.
-(2021) reimplements the MuJoCo simulator in JAX and also provides a PPO implementation as a baseline.
-Jax-LOB Frey et al.
-(2023) implements a vectorized limit order book as an RL environment that runs on the accelerator.
-Perhaps the most similar to our work is Mava Pretorius et al.
-(2021), which provides a MAPPO baseline, as well as integration with the Robot Warehouse environment.
-However, none of these libraries combine a range of JAX-based MARL environments with both value-based and actor-critic baselines.
-Broadly, no other work provides implementations of a wide range of hardware-accelerated MARL environments, while also implementing value-based and actor-critic baselines.
-Secondly, no other JAX simplification of SMAC exists.
-All other versions are either tied to the StarCraft II simulator or not hardware accelerated.
-7. Conclusion
-Hardware acceleration offers important opportunities for MARL research by lowering computational barriers, increasing the speed at which ideas can be iterated, and allowing for more thorough evaluation.
-We present JaxMARL, an open-source library of popular MARL environments and baseline algorithms implemented in JAX.
-We combine ease of use with hardware accelerator enabled efficiency to give significant speed-ups compared to traditional CPU-based implementations.
-Furthermore, by bringing together a wide range of MARL environments under one codebase, we have the potential to help alleviate issues with MARL’s evaluation standards.
-We hope that JaxMARL will help advance MARL by improving the ability of academic labs to conduct research with thorough, fast, and effective evaluations.
-8. Author Contributions
-This project is a large-scale effort spanning many labs and contributors.
-AR11footnotemark: 122footnotemark: 2 led the design of the JaxMARL API and interface the implementation of IPPO and MPE environments.
-BE11footnotemark: 122footnotemark: 2 led the design and implementation of the SMAX environments and IPPO evaluations.
-AR and BE also led the writing of this manuscript.
-MG11footnotemark: 122footnotemark: 2 led the implementation of the off-policy MARL algorithms, their evaluations., and the implementation of the Switch Riddle environment.
-JC11footnotemark: 1 led the implementation of the Hanabi environment and heavily assisted with benchmarking and verifying its performance.
-AL11footnotemark: 1 led the implementation of the Overcooked environments.
-GI11footnotemark: 1 led the implementation of the Multi-Agent Brax environments.
-TW11footnotemark: 1 led the implementation of the STORM environments.
-AK and AS worked on the STORM environments.
-CSW led the implementation of the Predator-Prey environment.
-CSW, SB, MS, MJ, and RL provided invaluable discussions for project planning and implementations across the project.
-SB helped initiate the project plan.
-MS worked on the Multi-Agent Brax environments.
-MJ worked on the Overcooked and Hanabi environments.
-RL assisted with the design of the API and testing infrastructure.
-SW, BL, NH, and TR provided invaluable feedback on the project, manuscript, and results.
-CL11footnotemark: 122footnotemark: 2 initiated the project and led the organizational and planning efforts, speed-based benchmarking, and Coin Game implementation.
-JF is the primary advisor for the project.
-Appendix A Further Details on Environments
-A.1. SMAX
-Observations in SMAX are structured similarly to SMAC.
-Each agent observes the health, previous action, position, weapon cooldown and unit type of all allies and enemies in its sight range.
-Like SMACv2Ellis et al.
-(2022), we use the sight and attack ranges as prescribed by StarCraft II rather than the fixed values used in SMAC.
-SMAX and SMAC have different returns.
-SMAC’s reward function, like SMAX’s, is split into two parts: one part for depleting enemy health, and another for winning the episode.
-However, in SMAC, the part which rewards depleting enemy health scales with the number of agents.
-This is most clearly demonstrated in 27m_vs_30m, where a random policy gets a return of around out of a maximum of because almost all the reward is for depleting enemy health or killing agents, rather than winning the episode.
-In SMAX, however, 50% of the total return is always for depleting enemy health, and 50% for winning.
-Unlike StarCraft II, where all actions happen in a randomised order in the game loop, some actions in SMAX are simultaneous, meaning draws are possible.
-In this case both teams get reward.
-Like SMAC, each environment step in SMAX consists of eight individual time ticks.
-SMAX uses a discrete action space, consisting of movement in the four cardinal directions, a stop action, and a shoot action per enemy.
-SMAX makes three notable simplifications of the StarCraft II dynamics to reduce complexity.
-First, zerg units do not regenerate health.
-This health regeneration is slow at health per second, and so likely has little impact on the game.
-Protoss units also do not have shields.
-Shields only recharge after 10 seconds out of combat, and therefore are unlikely to recharge during a single micromanagement task.
-Protoss units have additional health to compensate for their lost shields.
-Finally, the available unit types are reduced compared to SMAC.
-SMAX has no medivac, colossus or baneling units.
-Each of these unit types has special mechanics that were left out for the sake of simplicity.
-For the SMACv2 scenarios, the start positions are generated as in SMACv2, with the small difference that the ‘surrounded’ start positions now treat allies and enemies identically, rather than always spawning allies in the middle of the map.
-This symmetry guarantees that a 50% win rate is always achievable.
-Collisions are handled by moving agents to their desired location first and then pushing them out from one another.
-A.2. Coin Game
-Two agents, ‘red’ and ‘blue’, move in a wrap-around grid and collect red and blue coloured coins.
-When an agent collects any coin, the agent receives a reward of .
-However, when ‘red’ collects a blue coin, ‘blue’ receives a reward of and vice versa.
-Once a coin is collected, a new coin of the same colour appears at a random location within the grid.
-If a coin is collected by both agents simultaneously, the coin is duplicated and both agents collect it.
-Episodes are of a set length.
-A.3. Spatial-Temporal Representations of Matrix Games (STORM)
-This environment features directional agents within an 8x8 grid-world with a restricted field of view.
-Agents cannot move backwards or share the same location.
-Collisions are resolved by either giving priority to the stationary agent or randomly if both are moving.
-Agents collect two unique resources: cooperate and defect coins.
-Once an agent picks up any coin, the agent’s colour shifts, indicating its readiness to interact.
-The agents can then release an interact beam directly ahead; when this beam intersects with another ready agent, both are rewarded based on the specific matrix game payoff matrix.
-The agents’ coin collections determine their strategies.
-For instance, if an agent has 1 cooperate coin and 3 defect coins, there’s a 25% likelihood of the agent choosing to cooperate.
-After an interaction, the two agents involved are frozen for five steps, revealing their coin collections to surrounding agents.
-After five steps, they respawn in a new location, with their coin count set back to zero.
-Once an episode concludes, the coin placements are shuffled.
-This grid-based approach to matrix games can be adapted for n-player versions.
-While STORM is inspired by MeltingPot 2.0, there are noteworthy differences:
-Meltingpot uses pixel-based observations while we allow for direct grid access.
-Meltingpot’s grid size is typically 23x15, while ours is 8x8.
-Meltingpot features walls within its layout, ours does not.
-Our environment introduces stochasticity by shuffling the coin placements, which remain static in Meltingpot.
-Our agents begin with an empty coin inventory, making it easier for them to adopt pure cooperate or defect tactics, unlike in Meltingpot where they start with one of each coin.
-MeltingPot is implemented in Lua (Ierusalimschy, 2006) where as ours is a vectorized implementation in Jax.
-We deem the coin shuffling especially crucial because even large environments representing POMDPs, such as SMAC, can be solved without the need for memory if they lack sufficient randomness (Ellis et al., 2022).
-A.4. Hanabi
-There are a few details that differ between our Hanabi implementation and the original Hanabi Learning Environment (HLE).
-The most notable of these is how we choose to represent card knowledge information in the agents’ observation.
-In the HLE, card knowledge is observed as a colour/rank if there has been an explicit hint about a given card.
-As a separate feature, implicit card knowledge is represented as possible colours/ranks if there has not been an explicit hint that indicates a given card is not that colour/rank.
-We, on the other hand, combine implicit and explicit card knowledge, by only maintaining a representation of implicit card knowledge, which reduces to explicit card knowledge in the event an explicit hint is given about a card.
-This is because all possible colours/ranks are represented as 1s, whilst all ruled out colours/ranks are represented as 0s.
-By giving an explicit hint, all but one colour/rank are ruled out, leaving a one-hot encoding of the explicit card knowledge.
-We implement card knowledge this way, because knowledge updates are implemented via tensor calculus using JAX Numpy arrays of fixed shape and data type.
-Appendix B Value-Based MARL Methods and Implementation details
-Key features of our framework include parameter sharing, a recurrent neural network (RNN) for agents, an epsilon-greedy exploration strategy with linear decay, a uniform experience replay buffer, and the incorporation of Double Deep -Learning (DDQN) Van Hasselt et al.
-(2016) techniques to enhance training stability.
+Appendix D Value-Based MARL Methods and Implementation details
+Key features of our framework include parameter sharing, a recurrent neural network (RNN) for agents, an epsilon-greedy exploration strategy with linear decay, a uniform experience replay buffer, and the incorporation of Double Deep Q-Learning (DDQN) [64] techniques to enhance training stability.
+We stored the replay buffer in GPU memory using Flashbax [63].
 Unlike PyMARL, we use the Adam optimizer as the default optimization algorithm. Below is an introduction to common value-based MARL methods.
-IQL (Independent -Learners) is a straightforward adaptation of Deep -Learning to multi-agent scenarios.
-It features multiple -Learner agents that operate independently, optimizing their individual returns.
+IQL (Independent Q-Learners) is a straightforward adaptation of Deep Q-Learning to multi-agent scenarios.
+It features multiple Q-Learner agents that operate independently, optimizing their individual returns.
 This approach follows a decentralized learning and decentralized execution pipeline.
-VDN (Value Decomposition Networks) extends -Learning to multi-agent scenarios with a centralized-learning-decentralized-execution framework.
+VDN (Value Decomposition Networks) extends Q-Learning to multi-agent scenarios with a centralized-learning-decentralized-execution framework.
 Individual agents approximate their own action’s -Value, which is then summed during training to compute a jointed for the global state-action pair.
@@
 These hypernetworks are two-layered multi-layer perceptrons with ReLU non-linearity.
-Appendix C Training Results
-C.1. MABrax
-The performance of IPPO on HalfCheeta and Walker is reported in Figure 9, with hyperparameters reported in Table 4.
-C.2. MPE
-Performance of -Learning baselines in all the MPE scenarios are reported in Figure 10.
+Issues found when using Q-Learning in an end-to-end GPU setting.
+As discussed in the paper’s results section, PPO demonstrates a clear advantage over Q-Learning for our benchmarked environments, both in agent performance and training runtime.
+The speed differential is caused by the optimal sampling/replay ratio for Q-Learning methods becoming rapidly unbalanced as the number of parallel environments increases, which requires us to use fewer parallel environments than we use with PPO.
+PPO also has a major advantage over Q-Learning in that it does not use a replay buffer, which can occupy a significant amount of GPU memory.
+Secondly, our experiments empirically showed PPO to be more stable during training.
+A possible workaround is to increase the replay ratio by performing multiple update steps per training episode, which nevertheless affects computational efficiency.
+A better solution is to implement a distributed framework, separating the learning and sampling process, which is also out-of-scope for this work.
+Appendix E Speed Comparison
+The runs reported in Figures 3 and 5(c) were all run on the same system featuring two NVIDIA GeForce RTX 4090s (although only one was used for training), an Intel(R) Xeon(R) Silver 4316 CPU (20 cores with 40 threads), and 132 GB of RAM.
+We report the average environment steps per second over the entire RL training process, which for JaxMARL includes any compilation time.
+For Table 3, all results were collected on a single NVIDIA A100 GPU and AMD EPYC 7763 64-core processor.
+Environments were rolled out for 1000 sequential steps.
+In Figure 11 we repeat the analysis, reported in the main paper for QMIX, of JaxMARL’s ability to train multiple seeds in parallel for IPPO.
+Training this way allows training agents many thousands of times faster, with a 12500x speed up in the MPE simple spread environment.
+Appendix F Training & Correctness Results
+F.1 Overcooked
+We train IPPO, VDN and IQL agents in Overcooked and present their aggregate performance in Figure 12(a).
+IPPO performs better than the Q-Learning methods in inter-quartile mean and mean, in line with our more general findings.
+During training, we use the same shaped reward as stated in the original Overooked paper [10], which is added to the score of the game with a factor that is decayed from 1 to 0 during the first half of training.
+We don’t train MAPPO and QMIX for this task because, in Overcooked, agents can observe the entire state of the map.
+Therefore, there is no partial observability that can be improved through centralized training.
+We demonstrate correspondence by training an IPPO policy with JaxMARL on our implementation and evaluating the policy over 10 rollouts for both our Overcooked implementation and the original.
+Results are shown in Figure 14 with the similarity in performance demonstrating their equivalence.
+F.2 MABrax
+The performance of IPPO on ant_4x2, humanoid_9|8, hopper_3x1 and walker2d_2x3 is reported in Figure 15, with hyperparameters reported in Table 7.
+F.3 MPE
+Performance of -Learning baselines in all the MPE scenarios are reported in Figure 17 and Figure 18.
 The upper row represents cooperative scenarios, with results for all our -learning baselines reported.
 The bottom row refers to competitive scenarios, and results for IQL are divided by agent types.
-Hyperparameters are given in Table 10
-C.3. SMAX
-The performance of IPPO in SMAX versus MAPPO in SMAC is shown in Figure 11 while the performance of our -learning baselines is reported in Figure 12.
-We do not report them together because their hyperparameters were tuned over a different number of timesteps.
-Hyperparameters for IPPO and the -learning methods are given in Tables 6 and 10 respectively.
-Appendix D Hyperparameters
-| Value | Ant | HalfCheetah | Walker |
-| VF_COEF | 4.5 | 0.14 | 1.9 |
-| ENT_COEF | |||
-| LR | |||
-| NUM_ENVS | 64 | – | – |
-| NUM_STEPS | 300 | – | – |
-| TOTAL_TIMESTEPS | – | – | |
-| NUM_MINIBATCHES | 4 | – | – |
-| GAMMA | 0.99 | – | – |
-| GAE_LAMBDA | 1.0 | – | – |
-| CLIP_EPS | 0.2 | – | – |
-| MAX_GRAD_NORM | 0.5 | – | – |
-| ACTIVATION | tanh | – | – |
-| ANNEAL_LR | True | – | – |
-| Hyperparameter | Value |
-| LR | |
-| NUM_ENVS | |
-| NUM_STEPS | |
-| TOTAL_TIMESTEPS | |
-| UPDATE_EPOCHS | |
-| NUM_MINIBATCHES | |
-| GAMMA | |
-| GAE_LAMBDA | |
-| CLIP_EPS | |
-| ENT_COEF | |
-| VF_COEF | |
-| MAX_GRAD_NORM | |
-| ACTIVATION | tanh |
-| ANNEAL_LR | True |
-| Hyperparameter | Value |
-| LR | |
-| NUM_ENVS | |
-| NUM_STEPS | |
-| TOTAL_TIMESTEPS | |
-| UPDATE_EPOCHS | |
-| NUM_MINIBATCHES | |
-| GAMMA | |
-| GAE_LAMBDA | |
-| CLIP_EPS | |
-| SCALE_CLIP_EPS | False |
-| ENT_COEF | |
-| VF_COEF | |
-| MAX_GRAD_NORM | |
-| ACTIVATION | relu |
-| Hyperparameter | Value |
-| LR | |
-| NUM_ENVS | |
-| NUM_STEPS | |
-| TOTAL_TIMESTEPS | |
-| UPDATE_EPOCHS | |
-| NUM_MINIBATCHES | |
-| GAMMA | |
-| GAE_LAMBDA | |
-| CLIP_EPS | |
-| ENT_COEF | |
-| VF_COEF | |
-| MAX_GRAD_NORM | |
-| ACTIVATION | relu |
-| ANNEAL_LR | True |
-| NUM_FC_LAYERS | 2 |
-| LAYER_WIDTH | 512 |
-| Hyperparameter | Value |
-| LR | |
-| NUM_ENVS | |
-| NUM_STEPS | |
-| TOTAL_TIMESTEPS | |
-| UPDATE_EPOCHS | |
-| NUM_MINIBATCHES | |
-| GAMMA | |
-| GAE_LAMBDA | |
-| CLIP_EPS | |
-| ENT_COEF | |
-| VF_COEF | |
-| MAX_GRAD_NORM | |
-| ACTIVATION | tanh |
-| ANNEAL_LR | True |
-| NUM_EVALS |
-| Hyperparameter | Value |
-| NUM_ENVS | |
-| NUM_STEPS | |
-| BUFFER_SIZE | |
-| BUFFER_BATCH_SIZE | |
-| TOTAL_TIMESTEPS | |
-| AGENT_HIDDEN_DIM | |
-| AGENT_INIT_SCALE | |
-| EPSILON_START | |
-| EPSILON_FINISH | |
-| EPSILON_ANNEAL_TIME | |
-| MIXER_EMBEDDING_DIM* | |
-| MIXER_HYPERNET_HIDDEN_DIM* | |
-| MIXER_INIT_SCALE* | |
-| MAX_GRAD_NORM | |
-| TARGET_UPDATE_INTERVAL | |
-| LR | |
-| EPS_ADAM | |
-| WEIGHT_DECAY_ADAM | |
-| GAMMA | |
-| NUM_TEST_EPISODES | |
-| TEST_INTERVAL |
-(* Parameters specific to QMix.)
-| Hyperparameter | Value |
-| NUM_ENVS | |
-| NUM_STEPS | |
-| BUFFER_SIZE | |
-| BUFFER_BATCH_SIZE | |
-| TOTAL_TIMESTEPS | |
-| AGENT_HIDDEN_DIM | |
-| AGENT_INIT_SCALE | |
-| EPSILON_START | |
-| EPSILON_FINISH | |
-| EPSILON_ANNEAL_TIME | |
-| MIXER_EMBEDDING_DIM* | |
-| MIXER_HYPERNET_HIDDEN_DIM* | |
-| MIXER_INIT_SCALE* | |
-| MAX_GRAD_NORM | |
-| TARGET_UPDATE_INTERVAL | |
-| LR | |
-| EPS_ADAM | |
-| WEIGHT_DECAY_ADAM | |
-| GAMMA | |
-| NUM_TEST_EPISODES | |
-| TEST_INTERVAL |
-(* Parameters specific to QMix.)
+Hyperparameters are given in Table 8.
+F.4 SMAX
+The performance of different algorithms in SMAX versus MAPPO in SMAC is shown in Figure 19.
+Hyperparameters for IPPO and the -learning methods are given in Table 6 and Table 8 respectively.
+Some maps are significantly more difficult in SMAX, such as 10m_vs_11m, whereas some are much easier such as 3s_vs_5z.
+F.5 Hanabi
+Appendix G Hyperparameters
+| Parameter | MPE | SMAX | Hanabi | Overcooked |
+| PPO | ||||
+| # training timesteps | ||||
+| # parallel environments | 16 | 64 | 1024 | 64 |
+| # rollout steps | 128 | 128 | 256 | |
+| Adam learning rate | ||||
+| Anneal learning rate | True | True | True | True |
+| Update epochs | 2 | 4 | 4 | |
+| Minibatches per epoch | 2 | 4 | 16 | |
+| 0.99 | 0.99 | 0.99 | ||
+| 0.95 | 0.95 | 0.95 | ||
+| Clip range | 0.2 | 0.2 | 0.2 | |
+| Entropy coefficient | 0.0 | 0.01 | 0.01 | |
+| Value loss coefficient | 0.5 | 0.5 | 0.5 | |
+| Maximum gradient norm | 0.5 | 0.5 | 0.5 | |
+| Activation | tanh | relu | relu | relu |
+| Feed-forward network | ||||
+| Number of layers | 2 | - | 2 | 2 |
+| Fully-connected width | 64 | - | 512 | 64 |
+| Recurrent network | ||||
+| Hidden width | 128 | 128 | 128 | - |
+| Number of full-connected layers | 2 | 2 | 2 | - |
+| Fully-connected width | 128 | 128 | 128 | - |
+| Parameter | Ant | HalfCheetah | Walker |
+| PPO | |||
+| # training timesteps | |||
+| # parallel environments | 64 | ||
+| # rollout steps | 300 | ||
+| Adam learning rate | |||
+| Anneal learning rate | True | ||
+| Update Epochs | 4 | ||
+| Minibatches per epoch | 4 | ||
+| 0.99 | |||
+| 1.0 | |||
+| Clip range | 0.2 | ||
+| Entropy coefficient | |||
+| Value loss coefficient | 4.5 | 0.14 | 1.9 |
+| Maximum gradient norm | 0.5 | ||
+| Activation | tanh | ||
+| Feed-forward network | |||
+| Number of layers | 2 | ||
+| Fully-connected width | 64 |
+| Parameter | MPE | SMAX | Overcooked |
+| Q-Learning | |||
+| # training timesteps | |||
+| # parallel environments | 8 | 16 | 32 |
+| # rollout steps | 26 | 128 | 1 |
+| Adam learning rate | |||
+| Anneal learning rate | True | False | True |
+| Buffer size | |||
+| Buffer batch size | 32 | 32 | 128 |
+| start | 1.0 | 1.0 | 1.0 |
+| finish | 0.05 | 0.05 | 0.05 |
+| decay | 0.1 | 0.1 | 0.2 |
+| Hidden size | 64 | 512 | 64 |
+| Maximum gradient norm | 25 | 10 | 1 |
+| 1.0 | 1.0 | 1.0 | |
+| Update epochs | 1 | 8 | 4 |
+| Learning starts at timestep | |||
+| 0.9 | 0.99 | 0.99 | |
+| QMIX specific | |||
+| Mixed embedding width | 32 | 64 | - |
+| Mixer hypernet width | 128 | 256 | - |
+| Mixer initial scale | - |